### 16 Mar 2026 v1

# A new model for the quantum mechanics of the Hydrogen atom

Abstract
In this paper we introduce a new model for the quantum-mechanical system of the hydrogen atom.
We start with a four-dimensional Lorentzian quadratic space $(\mathcal{V},q)$ and let $C\subset\mathcal{V}$ be the corresponding cone.
The Hilbert space of our model $\boldsymbol{H}$ consists of $L^{2}$ functions on the cone and observables are represented by operators in the algebra $\mathcal{D}(C)$ of algebraic differential operators on $C$. We introduce a distinguished Schwartz subspace $\boldsymbol{H}^{\infty}$ of $\boldsymbol{H}$ that is naturally a $\mathcal{D}(C)$-module. The Schrödinger operator in our system is represented by a Schrödinger family of operators in $\mathcal{D}(C)$.
We compute the spectrum of the Schrödinger family in the Schwartz space $\boldsymbol{H}^{\infty}$ and show that it coincides with the spectrum in physics and solutions in $\boldsymbol{H}^{\infty}$ correspond to the usual solutions in physics.
The main differences with the standard model are:
  1. 1.
    We use the cone $C$ instead of $\mathbb{R}^{3}$ as our configuration space. As a result, the group of geometric symmetries of our configuration space is $O(q)\simeq O(3,1)$ rather than $O(3)\ltimes\mathbb{R}^{3}$ for $\mathbb{R}^{3}$.
  2. 2.
    We only use algebraic operators with no singularities.
  3. 3.
    We do not use any specific boundary conditions for solutions of our equations. They are all hidden in the Schwartz space $\boldsymbol{H}^{\infty}$.
1 Introduction
1.1 Overview
The hydrogen atom system is a fundamental well-studied example of a quantum-mechanical system which played a central role in the development of quantum mechanics, see e.g., [Wybourne, LandauLifshitz, MR2156403, Wulfman]. The common model for the spinless non-relativistic system is given by the Hilbert space $L^{2}(\mathbb{R}^{3})$, together with the densely defined (normalized) Schrödinger operator
$H_{phys}=-\Delta-\frac{\kappa}{r},$}
where $\kappa$ is a positive constant, $r$ the radial coordinate and $\Delta$ the Laplacian on $\mathbb{R}^{3}$. Solutions for the time-independent Schrödinger equation,
$H_{phys}\psi=E\psi$
determine the time-evolution of the system, where $E\in\mathbb{R}$ is the energy and $\psi$ is a function on $\mathbb{R}^{3}\setminus\{0\}$ satisfying certain boundary conditions.
The spectrum of ${H}_{phys}$, that is the collection of all admissible energies, is
$\operatorname{Spec}({H}_{phys})=\left\{-\frac{\kappa^{2}}{(2n)^{2}}|n\in\mathbb{N}\right\}\cup[0,\infty).$
To deal with degeneracy of the energy eigenspaces one looks for solutions of the Schrödinger equation that are simultaneously eigenfunctions for the total angular momentum operator $L^{2}$ and for its projection on the $z$-axis $L_{z}$. We will refer to these solutions as physical solutions, and denote the linear span of such solutions for a given $E\in\operatorname{Spec}(H_{phys})$ by $\operatorname{PhysSol}(E)$. These solution spaces lead to several layers of symmetries in which the largest symmetry group is $O(4,2)$. In physics, it is called the dynamical symmetry e.g., see [sym12081323]. The model we just described has several flaws.
  • The Schrödinger operator seems ad hoc and definitely not canonical.
  • The Schrödinger operator has a singularity at the origin and it involves the non-algebraic square root function via the radial coordinate.
  • The Schrödinger operator is not directly a part of the action of the Lie algebra $\mathfrak{o}(4,2)$ that arises from the above-mentioned $O(4,2)$ symmetry.
  • The boundary conditions that are used for the physical solutions do not have a conceptually clear motivation.
The purpose of this paper is to present an algebraic model for the quantum-mechanical system of the hydrogen atom that does not suffer from the above-mentioned flaws and recovers the known spectrum, the physical solutions and their symmetries.
We associate a model with every Lorentzian quadratic space $(\mathcal{V},q)$. To simplify the discussion here we will start with $\mathcal{V}=\mathbb{R}^{4}$ equipped with the Lorentzian quadratic form $q(x,y,z,w)=x^{2}+y^{2}+z^{2}-w^{2}$.
We denote the corresponding cone by
$C(\mathbb{R})=\{v\in\mathbb{R}^{4}|q(v)=0\},$
and its regular part $C(\mathbb{R})\setminus\{0\}$ by $C_{0}(\mathbb{R})$. The regular part of the cone splits as a union of two connected components, the upper half cone $C_{+}(\mathbb{R})$, and the lower half cone $C_{-}(\mathbb{R})$.
The cone carries a canonical invariant measure $|\omega|$, which allows us to define the first ingredient of our model which is the canonical Hilbert space, $\boldsymbol{H}=L^{2}({C}_{0}(\mathbb{R}),|\omega|)$ of square integrable functions on the cone.
The second ingredient is the algebra $\mathcal{D}(C)$ of algebraic differential operators on $C(\mathbb{R})$.
There is a natural action $\rho^{\infty}$ of the algebra $\mathcal{D}(C)$ on $C^{\infty}(C_{0}(\mathbb{R}))$ the space of smooth functions on $C_{0}(\mathbb{R})$. We construct a distinguished subspace $\boldsymbol{H}^{\infty}$ of $\boldsymbol{H}$ that we call the Schwartz space of $\boldsymbol{H}$. The subspace $\boldsymbol{H}^{\infty}$ is contained in $C^{\infty}(C_{0}(\mathbb{R}))$ and invariant under the action of the algebra $\mathcal{D}(C)$. Moreover this action $\rho$ is self-adjoint. The self-adjoint action $(\rho,\boldsymbol{H}^{\infty})$ of the algebra $\mathcal{D}(C)$ is the third ingredient of our model. We remark that the Schwartz subspace $\boldsymbol{H}^{\infty}$ encodes the fact that $\boldsymbol{H}$ carries a distinguished unitary representation of an orthogonal group $G(\mathcal{V})$ (that extends $O(\mathcal{V})$ and is isomorphic to $O(4,2)$).
The last ingredient of our model is the Schrödinger family, which is a one-parameter family of differential operators on the cone $E\longmapsto S(E)\in\mathcal{D}(C)$, parameterized by $E\in\mathbb{C}$. Explicitly,
$S(E)=H_{w}+\kappa+Ew,$
where $H_{w}$ is a certain second order differential operator on $C$. The Schrödinger family corresponds to the Schrödinger operator $H_{phys}$ in the standard model of the Hydrogen atom.
We define the spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}$ to be the collection of all $E\in\mathbb{C}$ for which the kernel of $S(E)$ in $(\boldsymbol{H}^{\infty})^{*}$ (the dual space of $\boldsymbol{H}^{\infty}$) is non-zero.
The Schwartz space $\boldsymbol{H}^{\infty}$ as a representation of $\mathcal{D}(C)$ has a canonical decomposition into invariant subspaces $\boldsymbol{H}^{\infty}=\boldsymbol{H}^{\infty}_{+}\oplus\boldsymbol{H}^{\infty}_{-}$, where $\boldsymbol{H}^{\infty}_{\pm}$ consists of those functions in $\boldsymbol{H}^{\infty}$ that are supported in the half cone $C_{\pm}(\mathbb{R})$. We prove that the spectrum of the Schrödinger family in the upper half cone $\boldsymbol{H}^{\infty}_{+}$ is equal to the spectrum of the Schrödinger operator $H_{phys}$. Moreover the physical solution spaces can be identified with solution spaces (kernels) of the Schrödinger family in $\boldsymbol{H}^{\infty}_{+}$. We also show that the spectrum of the Schrödinger family in the lower half cone $\boldsymbol{H}^{\infty}_{-}$ is equal to $(0,\infty)$. These solutions are not seen in the standard description of the system in physics. It is natural to ask for a possible physical meaning of these new solutions. We do not know the answer to this question.
Our model is completely algebraic and does not suffer from the above-mentioned flaws of the classical model.
Before giving a detailed account of our main results we will make two comments that put this work within a larger context and have mathematical implications that go beyond this work.
The boundary conditions that are used in the classical model do not arise in our model and they are implicitly hiding in the Schwartz space $\boldsymbol{H}^{\infty}$. This property of implicit boundary conditions through the canonical smooth subspace may have uses elsewhere. For example, when solving differential equations within a unitary representation of a reductive Lie group.
Another point that we discuss below is a group-theoretical and representation-theoretical description of the spectrum problem for the Hydrogen atom. We show how the Schrödinger family correspond to a one-parameter family of elements in the universal enveloping algebra of $SL_{2}(\mathbb{R})$. Moreover we show that the calculation of the spectrum of the Schrödinger family in $\boldsymbol{H}$ can be reduced to calculations completely within discrete series representations of $SL_{2}(\mathbb{R})$.
We finish this part of the introduction by describing the main observation that suggested to us the existence of our new model. We noticed two a priori unrelated situations in which there is an expected symmetry of $SO(n)$ but in practice the symmetry is of a larger orthogonal group.
The first instance is the usual model of the hydrogen atom system, in which there is a clear symmetry of $SO(3)$ and a larger hidden symmetry [^1] In fact there are different hidden symmetry groups within the dynamical symmetry that are fibers of a nice family, See [MR4460278, MR3827131]. of $O(4,2)$ that is called the dynamical symmetry.
The other instance is of the algebra of differential operators $\mathcal{D}(C)$ of the complex algebraic variety $C$ that is determined by a quadratic cone in $\mathbb{C}^{4}$. It is easy to see that first order elements in $\mathcal{D}(C)$ naturally contains $\mathfrak{so}(4,\mathbb{C})$. Much less obvious is the fact that the subspace of second order elements in $\mathcal{D}(C)$ contains a copy of $\mathfrak{so}(6,\mathbb{C})$. This was the starting point in our search for a new model for the hydrogen atom system which eventually lead to the results presented in this paper.
1.2 Structure of the paper and main results
We review the main results of the paper following the chronological order in which they appear in the rest of the paper.
1.2.1 Lorentzian quadratic space
A Lorentzian quadratic space is an $n$-dimensional real vector space $\mathcal{V}$ along with a Lorentzian quadratic form $q$ that is of signature (n1,1)(n-1,1). We will mostly deal with a Lorentzian quadratic space together with a fixed linear functional $w$ such that $q^{*}(w)=-1$, where $q^{*}$ is the dual form on the dual vector space. We will refer to this data as a pointed Lorentzian quadratic space and denote this in short by PLQS. With every PLQS we associate a corresponding model for the quantum-mechanical system of the $n$-dimensional hydrogen atom. The model described in the introduction above is the one associated with the PLQS with $\mathcal{V}=\mathbb{R}^{4}$, $q$ is the standard Lorentzian quadratic form, represented by the diagonal matrix $\operatorname{diag}(1,1,1,-1)$, and $w$ being the linear functional corresponding to the last coordinate.
Throughout this part of the introduction (Section 1.2) we assume that $(\mathcal{V},q,w)$ is a four-dimensional PLQS. This case corresponds to the case of the hydrogen atom system. We denote the complexification of $(\mathcal{V},q)$ by $(V,Q)$. The corresponding cone $C$ is the complex algebraic variety $\{v\in{V}|Q(v)=0\}$.
1.2.2 The algebra of differential operators on the cone
In Section 2 we study the algebra $\mathcal{D}(C)$ of differential operators on $C$. Later on we study its action on smooth functions on the corresponding real cone.
The algebra $\mathcal{D}(C)$ is filtered by degree of differential operators and each piece in the filtration has a grading according to homogeneity. This gives a decomposition
$\mathcal{D}(C)=\bigcup_{k\in\mathbb{N}_{0}}\left(\mathcal{D}^{k}(C)\right)=\bigcup_{k\in\mathbb{N}_{0}}\left(\oplus_{\ell\in\mathbb{Z}}\mathcal{D}^{k,\ell}(C)\right),$
where $\mathbb{N}_{0}=\mathbb{Z}_{\geq 0}$. It is known that the algebra $\mathcal{D}(C)$ is not generated by differential operators of degree one but it is generated by differential operators of degree two. Namely let
$\tilde{\mathfrak{s}}:=\oplus_{\ell+k=1,k\leq 2}\mathcal{D}^{k,\ell}(C)=\mathcal{D}^{0,1}(C)\oplus\mathcal{D}^{1,0}(C)\oplus\mathcal{D}^{2,-1}(C),$
it is clear that the $\tilde{\mathfrak{s}}$ is a Lie-algebra . The derived Lie-algebra $\mathfrak{s}:=[\tilde{\mathfrak{s}},\tilde{\mathfrak{s}}$], is isomorphic to $\mathfrak{o}(6,\mathbb{C})$ and moreover generates $\mathcal{D}(C)$.
The Lorentzian real form $(\mathcal{V},q)$ of $(V,Q)$ lead to a real form $\mathfrak{s}(\mathbb{R})$ of $\mathfrak{s}$. In section 8.2.2 we show that $\mathfrak{s}(\mathbb{R})$, the algebra of differential operators on the cone with real coefficients is isomorphic to $\mathfrak{o}(4,2)$.
1.2.3 The special morphism in $\mathcal{D}(C)$
The form $Q$ defines a canonical morphism $\Phi^{Q}$ of the vector space $\mathcal{D}(C)$. Explicitly, if in coordinates $Q$ is given by $\sum_{i,j=1}^{n}B_{ij}z_{i}z_{j}$, then for any $d\in\mathcal{D}(C)$,
$\Phi^{Q}(d)=\sum_{i,j=1}^{4}B_{ij}z_{i}dz_{j}.$
This is being discussed in Section 2.2.3. We describe the restriction of $\Phi^{Q}$ to the subspace $\mathcal{D}^{2,-1}(C)$ of $\mathfrak{s}$. We use the following result.
Proposition.
The morphism $\Phi^{Q}$ defines an isomorphism
$\Phi^{Q}|_{\mathcal{D}^{2,-1}(C)}:\mathcal{D}^{2,-1}(C)\xrightarrow{\sim}\mathcal{D}^{0,1}(C)=V^{*}.$
In fact we implicitly encountered the isomorphism $\Phi^{Q}|_{\mathcal{D}^{2,-1}(C)}$ before. Namely, the second order differential operator $H_{w}$ that appeared in the above-mentioned Schrödinger family $S(E)$ is defines using this isomorphism. More on this point can be found in sections 1.2.8 and 6 below.
In Appendix 8 we give a group theoretical characterization of $\Psi^{Q}$.
1.2.4 The dual pair symmetry
In Section 3.0.2 we show that $\mathfrak{s}(\mathbb{R})$ has a canonical dual pair, isomorphic to $(\mathfrak{so}(3),\mathfrak{sl}_{2}(\mathbb{R}))$ in $\mathfrak{so}(4,2)$. This explains how the PLQS give rise to a dual pair symmetry in $\mathcal{D}(C)$.
1.2.5 The Hilbert space
In section 4.1 we describe a canonical invariant measure $|\omega|$ on $C(\mathbb{R})$. Using the measure we attach to our PLQS a canonical Hilbert space $\boldsymbol{H}=L^{2}({C}(\mathbb{R}),|\omega|)$, consisting of square integrable functions on $C(\mathbb{R})$.
1.2.6 Dense actions
In Section 5.1 we recall the notion of a partial action of a *-algebra AA on a Hilbert space $\boldsymbol{H}$. A partial action of $A$ on $\boldsymbol{H}$ is given by an action $\rho:A\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\mathcal{W})$ of the *-algebra AA, on a linear subspace $\mathcal{W}$ of $\boldsymbol{H}$. The subspace $\mathcal{W}$ is called the domain of the action.
When $W$ is dense the partial action is called a dense action.
For a dense action there exists a well-defined adjoint partial action $\rho^{\dagger}:A\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\mathcal{W}^{\dagger})$.
We describe related notions of symmetric, self-adjoint and essentially self-adjoint dense actions of a *-algebra on a Hilbert space.
Our treatment is a natural generalization of the theory of unbounded operators on Hilbert spaces.
1.2.7 The self-adjoint action on the Schwartz space
There is a natural action $\rho^{\infty}$ of the algebra $\mathcal{D}(C)$ on the space $C^{\infty}(C_{0}(\mathbb{R}))$ of smooth functions on $C_{0}(\mathbb{R})$.
In Section 5.2 we show that $\rho^{\infty}$ restricts to a symmetric action on the domain $C^{\infty}_{c}(C_{0}(\mathbb{R}))$ of smooth compactly supported functions on $C_{0}(\mathbb{R})=C(\mathbb{R})\setminus\{0\}$.
Our main goal in Section 5 is to obtain a distinguished Schwartz subspace $\boldsymbol{H}^{\infty}$ of $\boldsymbol{H}$ that carries a self-adjoint action of the algebra $\mathcal{D}(C)$ extending the action on $C^{\infty}_{c}(C_{0}(\mathbb{R}))$. In order to do that in Section 5.3 we extend the group $O(q)\simeq O(3,1)$ to a group $G(\mathcal{V})$ isomorphic to $O(4,2)$. Then we show that the natural unitary representation of the group $O(q)$ on the canonical Hilbert space $\boldsymbol{H}$ can be extended to a distinguished unitary irreducible representation $\pi$ of $G(\mathcal{V})$ on the Hilbert space $\boldsymbol{H}.$
We define the Schwartz space $\boldsymbol{H}^{\infty}$ to be the canonical subspace of $\boldsymbol{H}$ consisting of smooth vectors for $\pi$. Using the inclusions
$C^{\infty}_{c}(C_{0}(\mathbb{R}))\subseteq\boldsymbol{H}^{\infty}\subseteq C^{\infty}(C_{0}(\mathbb{R}))$
and that $\boldsymbol{H}^{\infty}$ is stable under the action $\rho^{\infty}$ of the algebras $\mathcal{D}(C)$ we observe that $\boldsymbol{H}^{\infty}$ is a self-adjoint extension of the symmetric action on $C^{\infty}_{c}(C_{0}(\mathbb{R}))$.
1.2.8 The Schrödinger family
In Section 6 with every PLQS $(\mathcal{V},q,w)$ such that $q^{*}(w)=-1$, and every $\kappa\in\mathbb{R}$, we associate a one-parameter family of elements in the Lie algebra $\mathfrak{s}$, the Schrödinger family via
$S(q,w,\kappa;\lambda):=\Psi^{Q}\left({w}\right)+\kappa+\lambda{w}\in\mathfrak{s},\quad\lambda\in\mathbb{C}.$}
The parameter $\kappa$ plays the role of the strength of the Coulomb potential. For every $\lambda\in\mathbb{C}$, the operator $S(q,{w},\kappa;\lambda)$ belongs to the universal enveloping algebra of a canonical subalgebra $\ell$ of $\mathfrak{s}$, where $\ell\simeq\mathfrak{sl}_{2}(\mathbb{C})$. This gives rise to a one-parameter family in the universal enveloping algebra of $\mathfrak{sl}_{2}(\mathbb{C})$ (that we call the abstract Schrödinger family).
1.2.9 The model
Altogether, starting with a four-dimensional PLQS and $\kappa\in\mathbb{R}$, the corresponding model for the quantum-mechanical system of the hydrogen atom consists of
  1. 1.
    The canonical Hilbert space, $\boldsymbol{H}=L^{2}({C}_{0}(\mathbb{R})),$
  2. 2.
    The *-algebra of differential operators on the cone, $\mathcal{D}(C)$,
  3. 3.
    The distinguished self-adjoint action $(\rho,\boldsymbol{H}^{\infty})$, of the algebra $\mathcal{D}(C)$ on $\boldsymbol{H}$,
  4. 4.
    The Schrödinger family
    $S(q,w,\kappa;\lambda)=\Psi^{Q}\left({w}\right)+\kappa+\lambda{w}\in\mathcal{D}(C).$
1.2.10 The spectrum
In section 7 we define the spectrum of a family of operators on a topological vector space.
Under the action of the connected component $G(\mathcal{V})_{0}\simeq SO_{0}(4,2)$ the Hilbert space $\boldsymbol{H}$ has a canonical decomposition $\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}$ into irreducible sub-representations. The subspace $\boldsymbol{H}_{+}$ consists of functions in $\boldsymbol{H}$ that are supported in the upper half cone ans similarly $\boldsymbol{H}_{-}$ consists of functions that are supported in the lower half cone . We set $\boldsymbol{H}^{\infty}_{\pm}:=\boldsymbol{H}^{\infty}\cap\boldsymbol{H}_{\pm}$. We prove that the spectrum of the Schrödinger family in $\boldsymbol{H}_{+}^{\infty}$ is equal to the spectrum of the Schrödinger operator in physics. Below is the exact statement.
Theorem.
Let $(\mathcal{V},q,w)$ be a $4$-dimensional PLQS and let $\boldsymbol{H}$ be the corresponding canonical Hilbert space. For any $\kappa>0$,
$\operatorname{Spec}_{\boldsymbol{H}_{\pm}^{\infty}}\{S^{Ab}(q,w,\kappa;\lambda)|\lambda\in\mathbb{C}\}=\begin{cases}\left\{-\frac{\kappa^{2}}{4n^{2}}|n\in\mathbb{N}\right\}\cup[0,\infty),&+\\ (0,\infty),&-.\end{cases}$
In particular the spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}_{+}$ is equal to $\operatorname{Spec}(H_{phys})$ the spectrum in the standard model in physics. We identify the solution spaces for the Schrödinger family in $\boldsymbol{H}_{+}^{\infty}$ with the physical solutions. The spectrum in the full smooth space $\boldsymbol{H}^{\infty}$ has more solutions for $\lambda>0$ than in the classical model in physics.
1.2.11 Concluding remarks
Our suggested model does not suffer from the flaws of the classical model. It is defined with respect to a four-dimensional PLQS and $\kappa\in\mathbb{R}$. Up to a natural isomorphism, the model does not depend on the PLQS and in that sense it is canonical.
The main differences with the standard model are
  1. 1.
    We use the cone $C$ instead of $\mathbb{R}^{3}$ as our configuration space. As a result, the group of geometric symmetries of our configuration space is $O(q)\simeq O(3,1)$ rather than $O(3)\ltimes\mathbb{R}^{3}$ for $\mathbb{R}^{3}$.
  2. 2.
    We only use algebraic operators with no singularities and in particular the Schrödinger family belongs to $\tilde{\mathfrak{s}}\subset\mathcal{D}(C)$ and lead to a reductive dual pair symmetry of the type of $(\mathfrak{so}(3),\mathfrak{sl}_{2}(\mathbb{R}))$ in $\mathfrak{so}(4,2)$.
  3. 3.
    We do not use any specific boundary conditions for solutions of our equations. The distinguished Schwartz space $\boldsymbol{H}^{\infty}$ encodes them in an indirect way.
We compute the spectrum of the Schrödinger family in the Schwartz space $\boldsymbol{H}^{\infty}$ and show that it coincides with the spectrum in physics and solutions in $\boldsymbol{H}_{+}^{\infty}$ correspond to the usual solutions in physics.
This research was supported by the ISRAEL SCIENCE FOUNDATION (grant No. 1040/22).
2 The algebra of differential operators on a cone
Let $(V,Q)$ be a finite-dimensional complex nondegenerate quadratic and $C\subset V$ be the associated cone. In this section we discuss some known and some less known facts about the algebra $\mathcal{D}(C)$ of differential operators on $C$. We will specify and study a canonical Lie-sub-algebra $\mathfrak{s}$ of $\mathcal{D}(C)$ that generates it. We will also study related real structures of $\mathcal{D}(C)$.
2.1 Differential operators
2.1.1 Differential operators on linear spaces
In this work we will consider complex (and occasionally also real) affine algebraic varieties, which for our purpose can be given as the solution set of a system of polynomial equations in a complex affine space. For such a variety $X\subset\mathbb{C}^{n}$, its ring of polynomial functions $\mathcal{O}_{X}$ is a quotient of $\mathbb{C}[z_{1},z_{2},...,x_{n}]$. Grothendieck defined a differential operator of degree less or equal to $k\in\mathbb{N}_{0}$ on $X$ to be a linear operator $D\in\operatorname{End}_{\mathbb{C}}(\mathcal{O}_{X})$, for which
${f_{0}},[{f_{1}},...,[{f_{k}},D]]...]=0,$
for any $f_{0},f_{1},...,f_{k}\in\mathcal{O}_{X}$, and where we treat $\mathcal{O}_{X}$ as a subalgebra of $\operatorname{End}_{\mathbb{C}}(\mathcal{O}_{X})$. The collections $\mathcal{D}^{k}(X)$ of all differential operators of degree less or equal to $k$ on $X$ defines an increasing exhaustive filtration of $\mathcal{D}(X)$, the algebra of differential operators on $X$.
When $X$ is smooth, $\mathcal{D}(X)$ is generated as an algebra by $\mathcal{O}_{X}$ and $\operatorname{Der}(\mathcal{O}_{X})=\{D\in\mathcal{D}^{1}(X)|D(1)=0\}$. In general, beyond the smooth case, $\mathcal{D}(X)$ may not even be Noetherian, e.g., see the case of the cubic cone [cubiccone].
Consider the case in which the algebraic variety $X$ is a finite-dimensional complex vector space $V$. In this case we have an embedding $V^{*}\hookrightarrow\mathcal{O}_{X}\hookrightarrow\mathcal{D}^{0}(X)$ sending $\varphi$ to multiplication by $\varphi$ as an element in $D^{0}(V)$. In addition, we have an embedding of $V$ in $D^{1}(V)$ sending a vector $v$ to the corresponding directional derivative $\partial_{v}$. We note that for any $\varphi\in V^{*}$ and $v\in V$,
$\partial_{v},{\varphi}]=\partial_{v}(\varphi).$
The algebra $\mathcal{D}(V)$ of differential operators on $V$ is generated by the images of $V$ and $V^{*}$ in $\mathcal{D}(V)$. This algebra is called the Weyl algebra of $V$.
2.1.2 Differential operators on cones
From now on we will assume that $(V,Q)$ is a complex nondegenerate quadratic space with $\operatorname{dim}(V)=n>2$. We denote the associated symmetric bilinear form by $B_{Q}$ or by $B$, and the corresponding cone by
$C=C_{Q}=\{v\in V|Q(v)=0\}.$
In contrast to the smooth case, $\mathcal{D}(C)$ is not generated by $\mathcal{D}^{1}(C)$. However, $\mathcal{D}(C)$ is generated by $\mathcal{D}^{2}(C)$, and moreover it is isomorphic to a quotient of the universal enveloping algebra of $\mathfrak{so}(n+2,\mathbb{C})$.
We denote by $I=I_{Q}$ the ideal of $\mathcal{O}_{V}$ that is generated by $Q\in\mathcal{O}_{V}$.
The ring $\mathcal{O}_{C}$ of polynomial functions on $C$ is given by $\mathcal{O}_{V}/I$.
We say that a differential operator $D\in\mathcal{D}(V)$ is a lift of a differential operator $D^{\prime}\in\mathcal{D}(C)$ (or equivalently $D^{\prime}$ is a restriction of $D$) if
$D^{\prime}(f|_{C})=D(f)|_{C},\quad\forall f\in\mathcal{O}_{V}.$
A differential operator $D\in\mathcal{D}(V)$ has a restriction to a differential operator on $C$ if and only if $D(I)\subset I$. We define the subalgebra of $\mathcal{D}(V)$ consisting of all restrictable operators by
$\mathcal{D}(V)_{I}=\{D\in\mathcal{D}(V)|D(I)\subset I\}.$
The following result which follows from Proposition 1.6 of [SSP], explains how a differential operator on the cone can be described by differential operators on the ambient vector space.
Proposition.
The natural map $\mathcal{D}(V)_{I}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ is onto and having $I\mathcal{D}(V)$ as its kernel.
From now on we identify $\mathcal{D}(C)$ with $\mathcal{D}(V)_{I}/I\mathcal{D}(V)$.
2.2 Fine structure of the algebra $\mathcal{D}(C)$
2.2.1 The $\mathbb{Z}$-grading
The group $\mathbb{C}^{\times}$ acts via rescaling on $\mathcal{O}_{V}$. This induces an action on $\mathcal{D}(V)$ and under this action for each $k\in\mathbb{N}_{0}$, $\mathcal{D}^{k}(V)$ is invariant and hence carries a $\mathbb{Z}$-grading:
$\mathcal{D}^{k}(V)=\oplus_{\ell\in\mathbb{Z}}\mathcal{D}^{k,\ell}(V),$
with $\mathcal{D}^{k,\ell}(V)=\left\{d\in\mathcal{D}^{k}(V)|\alpha\cdot d=\alpha^{\ell}d,\forall\alpha\in\mathbb{C}^{\times}\right\}$. Moreover, since $C$ is stable under the action of $\mathbb{C}^{\times}$, the same holds for $\mathcal{D}(C)$:
$\mathcal{D}^{k}(C)=\oplus_{\ell\in\mathbb{Z}}\mathcal{D}^{k,\ell}(C).$
For any $k\in\mathbb{N}_{0}$ and $\ell\in\mathbb{Z}$, we set $\mathcal{D}^{k}(V)_{I}:=\mathcal{D}(V)_{I}\cap\mathcal{D}^{k}(V)$ and $\mathcal{D}^{k,\ell}(V)_{I}:=\mathcal{D}(V)_{I}\cap\mathcal{D}^{k,\ell}(V)$. Since $\mathcal{D}^{k}(V)_{I}$ is stable under the algebraic action of $\mathbb{C}^{\times}$ we must have
$\mathcal{D}^{k}(V)_{I}=\oplus_{\ell\in\mathbb{Z}}\mathcal{D}^{k,\ell}(V)_{I}.$
2.2.2 The canonical generating Lie subalgebra
Long ago it was discovered by I.M. Gelfand, S.I. Gelfand, and the first author that $\mathcal{D}(C)$ is generated by the Lie algebra
$\tilde{\mathfrak{s}}:=\oplus_{\ell+k=1,k\leq 2}\mathcal{D}^{k,\ell}(C)=\mathcal{D}^{0,1}(C)\oplus\mathcal{D}^{1,0}(C)\oplus\mathcal{D}^{2,-1}(C).$
We let $\mathfrak{s}$ be the derived Lie algebra $\tilde{\mathfrak{s}},\tilde{\mathfrak{s}}$]. We will call the Lie algebra $\mathfrak{s}$ the quantum dynamical symmetry. It can be verified that $\tilde{\mathfrak{s}}=\mathfrak{s}\oplus\mathbb{C}$.
Theorem (Compare with Prop. 3.8 and Cor. 4.4 of [LevasseurAndStafford2017]).
The complex Lie algebra $\mathfrak{s}$ is isomorphic to $\mathfrak{so}(n+2,\mathbb{C})$ and it generates $\mathcal{D}(C)$ as a unital complex algebra.
For an explicit embedding of $\mathfrak{so}(n+2,\mathbb{C})$ in $\mathcal{D}(C)$ see section 9 of [LevasseurAndStafford2017] and Theorem 8.2.1. Further details can be found in [Levasseur1986, LevasseurSmithStafford1988, Goncharov1982] and section 1.1 of [KobayashiMano2011].
Remark.
We note that the Euler operator $\sum_{i}x_{i}\partial_{x_{i}}$ belongs to $\tilde{\mathfrak{s}}$ but not to $\mathfrak{s}$}. The shifted Euler operator $h:=2\sum_{i}x_{i}\partial_{x_{i}}+(n-2)$ does belong to $\mathfrak{s}$. The scalar shift (n2)(n-2) arises from the natural action on half-forms.
2.2.3 The isomorphism $\mathcal{D}^{0,1}(C)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}^{2,-1}(C)$
In this section we show that the form $Q$ induces a canonical endomorphism $\Phi^{Q}$ of the vector space $\mathcal{D}(C)$. We focus on its restriction to the subspace $\mathcal{D}^{2,-1}(C)$ of $\mathfrak{s}$ which will be needed later on when we define the Schrödinger family.
We let $B_{Q}$ be the bilinear symmetric form associated to $Q$. Recall that coordinate functions $\{z_{i}\}$ on $V$ belong to $\mathcal{O}_{C}$, and using the embedding $\mathcal{O}_{C}\hookrightarrow D(C)$, we can think of each $z_{i}$ as a differential operator on the cone.
Proposition.
There exists a linear morphism $\Phi^{Q}:\mathcal{D}(C)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ such that if the form $Q$ is given in coordinates by $\sum_{i,j=1}^{n}B_{ij}z_{i}z_{j}$, then for any $d\in\mathcal{D}(C)$,
$\Phi^{Q}(d)=\sum_{i,j=1}^{n}B_{ij}z_{i}dz_{j}.$
Moreover $\Phi^{Q}$ is $O(Q)$-equivaraint.
The proof is by a straightforward calculation.
Restriction of $\Phi^{Q}$ to $\mathcal{D}^{2,-1}(C)$ is an isomorphism onto $\mathcal{D}^{0,1}(C)$. Before showing this we shall define a map $\Psi^{Q}:\mathcal{D}^{0,1}(C)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}^{2,-1}(C)$ that will be proportional to the mentioned inverse. For technical reasons it is more convenient to work with this un-normalized version of the mentioned inverse.
We set $n=\operatorname{dim(V)}$ and recall that $h=2\sum_{i}x_{i}\partial_{x_{i}}+(n-2)\in\mathfrak{s}$. We let $\Box_{Q}$ be the canonical second-order differential operator on $V$ corresponding to the symmetric bilinear form $B_{Q}$. For example, if in coordinates $Q(v)=\sum_{i=1}^{n}A_{i}z_{i}^{2}$ then $\Box_{Q}=\sum_{i=1}^{n}A_{i}^{-1}(\partial_{z_{i}})^{2}$.
The quadratic form $Q$ induces an isomorphism $\tau:V\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow V^{*}$. In this way we can assign with any linear functional $w\in V^{*}=\mathcal{D}^{0,1}(C)$ a unique vector $v_{w}\in V$. We let $\partial_{w}^{Q}$ be the directional derivative corresponding to $v_{w}$.
We define a linear map $\Psi^{Q}:V^{*}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}^{2,-1}(V)$ by
$\Psi^{Q}(w)=w\Box_{Q}-h\partial_{w}^{Q},\quad\forall w\in V^{*}.$
Lemma.
The linear map $\Psi^{Q}$ is an isomorphism onto its image. The image is equal to $\mathcal{D}^{2,-1}(V)_{I}$, and
$\Phi^{Q}\circ\Psi^{Q}=(2-n)\mathbb{I}_{V^{*}}.$
Proof.
By direct calculations we see that $\Psi^{Q}(V^{*})\subseteq\mathcal{D}^{2,-1}(V)_{I}$ and for every $w\in V^{*}$, $\Phi^{Q}\circ\Psi^{Q}(w)=(2-n)w$. In addition it is easy to show (and can be deduced from [Goncharov1982]) that $\operatorname{dim}(\mathcal{D}^{2,-1}(V)_{I})\leq\operatorname{dim}(V)$. ∎
We recall that $V^{*}\simeq\mathcal{D}^{0,1}(C)$ and $\mathcal{D}^{2,-1}(V)_{I}\simeq\mathcal{D}^{2,-1}(C)$. By abuse of notation we denote the corresponding isomorphism $\mathcal{D}^{0,1}(C)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}^{2,-1}(C)$ also by $\Psi^{Q}$.
In Appendix 8.3 we relate $\Psi^{Q}$ to Jacobson-Morozov theorem.
2.2.4 The symmetry group
The group $G_{1}=O(Q)\times\mathbb{C}^{\times}$ naturally acts on $V$ in a way that preserves $C$. As a result $G_{1}$ acts on $\mathcal{D}(V)$ and $\mathcal{D}(C)$. The spaces $\tilde{\mathfrak{s}}$ and $\mathfrak{s}$ are $G_{1}$ stable. Since $\mathcal{D}^{0,1}(V)_{I}$, $\mathcal{D}^{1,0}(V)_{I}$, $\mathcal{D}^{2,-1}(V)_{I}$ trivially intersect $I\mathcal{D}(V)$, the map from Proposition 2.1.2 induces a canonical $G$-equivariant isomorphism $\mu$ of complex Lie algebras from $\mathfrak{s}$ onto its image
$\mu(\tilde{\mathfrak{s}}):=\mathcal{D}^{0,1}(V)\oplus\mathcal{D}^{1,0}(V)_{I}\oplus\mathcal{D}^{2,-1}(V)_{I}.$
We will freely use the isomorphism $\mu$ throughout the text without explicitly mentioning it.
In addition, one can verify the following facts that we will use later on.
Proposition.
As representations of $G_{1}$ we have the following relations:
  1. 1.
    $\mathcal{D}^{0,1}(C)\cong\mathcal{D}^{0,1}(V)_{I}=V^{*}$.
  2. 2.
    $\mathcal{D}^{1,0}(C)\cong\mathcal{D}^{1,0}(V)_{I}=\mathfrak{o}(Q)\oplus\mathbb{C}1\oplus\mathbb{C}h$.
  3. 3.
    $\mathcal{D}^{2,-1}(C)\cong\mathcal{D}^{2,-1}(V)_{I}\cong V$.
  4. 4.
    $\mathcal{D}^{0,1}(C),\mathcal{D}^{2,-1}(C)]\cong\mathfrak{o}(Q)\oplus\mathbb{C}h$
In particular the decomposition of $\mathfrak{s}$ into irreducibles of $G_{1}$ is given by
$\mathfrak{s}\cong V\oplus\mathfrak{o}(Q)\oplus\mathbb{C}\oplus V^{*}.$
We remark that a description of $\mathcal{D}(C)$ is essentially given in [Goncharov1982]. We will also give an explicit description in the following section.
2.2.5 A spanning set for $\tilde{\mathfrak{s}}$
In this section we specify a spanning set for $\tilde{\mathfrak{s}}$ that can be used to prove several results that are stated in the following sections.
Since $\Psi^{Q}:\mathcal{D}^{0,1}(C)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}^{2,-1}(C)$, which was introduced in section 2.2.3, is an isomorphism, clearly
$\mathcal{D}^{0,1}(C)\cong\mathcal{D}^{0,1}(V)=\{\tau(v)|v\in V\},$
$\mathcal{D}^{2,-1}(C)\cong\mathcal{D}^{2,-1}(V)_{I}=\{\psi^{Q}(\tau(v))|v\in V\}.$
For $u,v\in V$ we define,
$L_{u,v}:=\tau(u)\circ\partial_{v}-\tau(v)\circ\partial_{u}\in\mathcal{D}^{1}(V),$
where $\partial_{v}$ is the directional derivative corresponding to $v$ and $\tau:V\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow V^{*}$ is the isomorphism that the quadratic form $Q$ induces. Note that $L_{u,v}\neq 0$ if and only if $\{u,v\}$ is linearly independent. By direct calculations we can show that
$\mathcal{D}^{1,0}(C)\cong\mathcal{D}^{1,0}(V)_{I}=\operatorname{span}_{\mathbb{C}}\{L_{u,v}|u,v\in V\}\oplus\mathbb{C}h\oplus\mathbb{C}.$
We remark that $\operatorname{span}_{\mathbb{C}}\{L_{u,v}|u,v\in V\}$ is canonically identified with the Lie algebras $\mathfrak{o}(Q)$.
Since $\tilde{\mathfrak{s}}=\mathcal{D}^{0,1}(C)\oplus\mathcal{D}^{1,0}(C)\oplus\mathcal{D}^{2,-1}(C)$ we have specified a generating set for $\tilde{\mathfrak{s}}$. The commutation relations of vectors in the specified spanning set are easy to calculate. They are given in Appendix 8.1.
2.2.6 Compatibility with real forms
The results presented in this section (Section 2) for complex quadratic spaces hold with appropriate changes to their real forms. Here we would like to point out two facts.
[…]
We remark that a real form of a complex vector space $V$ (or a quadratic space or a Lie algebra) is the same thing as an antiholomorphic involution $\sigma$ of $V$.
3 Quadratic spaces and reductive dual pairs
In this section we explain how a pointed real quadratic space gives rise to a reductive dual pair in $\mathfrak{s}$. This will play an important role for the Schrödinger family that is introduced in Section 6.1.4.
3.0.1 Pointed Lorentzian quadratic spaces
Our convention for a Lorentzian quadratic form on an $n$-dimensional vector space $\mathcal{V}$ is such that the signature of $q$ is (n1,1)(n-1,1).
Definition.
A pointed Lorentzian quadratic space (PLQS) is a triplet $(\mathcal{V},q,w)$ such that $(\mathcal{V},q)$ is a finite-dimensional (real) Lorentzian quadratic space and $w$ is a linear functional on $V$ of norm 1-1. That is, under the dual form $q^{*}$ on $\mathcal{V}^{*}$, $q^{*}(w)=-1$.
Remark.
The form $q$ induces a canonical isomorphism $\mathcal{V}\cong\mathcal{V}^{*}$ and in this way the dual space also carries a structure of a PLQS.
Throughout the text we will also work with the complexification $(V,Q)$ of $(\mathcal{V},q)$.
3.0.2 The reductive dual pair symmetry of a PLQS
In this section with a PLQS we associate a reductive dual pair in the quantum dynamical symmetry $\mathfrak{s}$.
Throughout this paragraph, we let $(\mathcal{V},q,w)$ be an $n$-dimensional PLQS with $n>2$.
Howe introduced the notion of a reductive dual pair, see e.g., [MR789290]. Given a reductive Lie algebra $\mathfrak{g}$ and a pair of reductive Lie subalgebras $(\mathfrak{k},\mathfrak{l})$ of $\mathfrak{g}$, the pair is called a reductive dual pair if $\mathfrak{k}$ is equal to the centralizer of $\mathfrak{l}$ in $\mathfrak{g}$ and $\mathfrak{l}$ is equal to the centralizer of $\mathfrak{k}$ in $\mathfrak{g}$.
The orthogonal group $O(Q)$ acts on the dual space $V^{*}$. The stabilizer ${K}_{w}:=\operatorname{stab}_{O(Q)}(w)$ is isomorphic to $O(n-1,\mathbb{C})$. We let $\mathfrak{k}_{w}\subset\mathcal{D}^{1,0}(C)$ be the complex Lie algebra of $K_{w}$. We define $\ell_{w}:=\operatorname{Cent}_{\mathfrak{s}}(\mathfrak{k}_{w})$.
We also note that the real Lie algebra $O(q)$ is given by $\operatorname{span}_{\mathbb{R}}\{L_{u,v}|u,v\in\mathcal{V}\}$ and isomorphic to $\mathfrak{so}(n-1,1)$. By Proposition 8.2.2 the real form $\mathfrak{s}(\mathbb{R})$ is isomorphic to $\mathfrak{so}(n,2)$. We define $\mathfrak{k}_{w}(\mathbb{R}):=\mathfrak{k}_{w}\cap\mathfrak{s}(\mathbb{R})$ and $\mathfrak{l}_{w}(\mathbb{R}):=\mathfrak{l}_{w}\cap\mathfrak{s}(\mathbb{R})$.
Proposition.
Keeping the above-mentioned notation.
  1. 1.
    The pair $(\mathfrak{k}_{w},\mathfrak{l}_{w})$ is a reductive dual pair in $\mathfrak{s}$. It is isomorphic to $(\mathfrak{o}(n-1,\mathbb{C}),\mathfrak{sl}_{2}(\mathbb{C}))$ in $\mathfrak{so}(n+2,\mathbb{C})$.
  2. 2.
    The pair $(\mathfrak{k}_{w}(\mathbb{R}),\l _{w}(\mathbb{R}))$ is a real reductive dual pair in $\mathfrak{s}(\mathbb{R})$ that is isomorphic to $(\mathfrak{so}(n-1),\mathfrak{sl}_{2}(\mathbb{R}))$ in $\mathfrak{so}(n,2)$.
The proof is by direct calculations and is given in Appendix 8.4
4 The canonical Hilbert space $\boldsymbol{H}$.
In this section, starting with an $n$-dimensional PLQS $(\mathcal{V},q,w)$ we construct a canonical Hilbert space $\boldsymbol{H}$.
Throughout this section we will assume that we are given an $n$-dimensional PLQS $(\mathcal{V},q,w)$ with $n$ being an even integer such that $n>2$.
4.1 The canonical Hilbert space
In this section, starting with a PLQS $(\mathcal{V},q,w)$ we introduce a canonical positive density $|\omega|$ on the corresponding real cone $C_{0}(\mathbb{R})$. We define the canonical Hilbert space $\boldsymbol{H}$ of our PLQS to be the corresponding $L^{2}$ space.
4.1.1 An orientation of a complex quadratic space
Definition.
Let $(V,Q)$ be a complex $n$-dimensional quadratic space. An orientation of $(V,Q)$ is a top form $\Omega$ on $V$ such that for any set of $n$ vectors $\{v_{1},v_{2},...,v_{n}\}\subset V$,
$\left(\Omega(v_{1},v_{2},...,v_{n})\right)^{2}=\det\left(G\right),$
where $G$ is the Gram matrix given by $G_{i,j}=B(v_{i},v_{j})$, and $B$ is the bilinear symmetric form corresponding to $Q$.
An oriented quadratic space is a quadratic space with a choice of orientation.
It is easy to show that an orientation exists and is unique up to a sign.
Example.
If $Q$ is the quadratic form on $\mathbb{C}^{n}$ that is given in coordinates by $Q=z_{1}^{2}+z_{2}^{2}+...+z_{n}^{2}$, the two orientations of $(\mathbb{C}^{n},Q)$ are given by
$\pm dz_{1}\wedge dz_{2}\wedge...\wedge dz_{n}.$
4.1.2 An orientation of a real quadratic space
Definition.
An orientation of a real quadratic space $(\mathcal{V},q)$ is an orientation $\Omega$ of its complexification $(V,Q)$.
An oriented real quadratic space is a quadratic space with a choice of orientation (for its complexification).
Example.
If $q$ is the quadratic form on $\mathbb{R}^{n}$ with signature $(\ell,m)$ that is given in coordinates by $Q=x_{1}^{2}+...+x_{\ell}^{2}-x_{\ell+1}^{2}-...-x_{\ell+m}^{2}$. The two orientations of $(\mathbb{R}^{n},q)$ are given by
$\pm(i)^{m}dx_{1}\wedge...\wedge dx_{n}.$
4.1.3 The canonical top form on $C_{0}$
Let $(V,Q,\Omega)$ be an oriented complex quadratic space.
We denote by $C_{0}$ the variety of smooth points of the corresponding complex cone $C$. Note that $C_{0}:=\{v\in V|Q(v)=0\}\setminus\{0\}$.
Claim.
There is a unique algebraic differential (n1)(n-1)-form $\omega$ on $C_{0}$ characterized by the following property. For any extension $\widetilde{\omega}$ of $\omega$ to a rational (n1)(n-1)-form on $V$, the rational function $f$ on $V$ that is defined by $dQ\wedge\widetilde{\omega}=f\Omega$ is equal to $1$ on $C_{0}$.
The form $\omega$ can be constructed locally as follows. For every $p\in C_{0}$, there is an open neighborhood $U_{p}$ of $p$ in $V$ and an (n1)(n-1)-form $\eta_{p}$ on $U_{p}$ such that $dQ\wedge\eta_{p}=\Omega$ on $U_{p}$. On $U_{p}\cap C_{0}$, $\omega|_{U_{p}\cap C_{0}}=\eta_{p}|_{U_{p}\cap C_{0}}$.
Hence an orientation $\Omega$ on $(V,Q)$ gives rise to a canonical top form $\omega$ on $C_{0}$. Clearly $\omega$ is $SO(Q)$-invariant.
4.1.4 The canonical density on a Lorentzian space
In this section we show that a Lorentzian quadratic space caries a canonical density.
We let $(\mathcal{V},q)$ be a real Lorentzian oriented quadratic space with complexification $(V,Q)$ and orientation $\Omega$. The real cone $C_{0}(\mathbb{R}):=C_{0}\cap\mathcal{V}$ is a smooth manifold. Restriction of the canonical form $\omega$ to $C_{0}(\mathbb{R})$ gives a complex-valued smooth (n1)(n-1)-form.
Taking the absolute value defines a canonical positive density $|\omega|$ on $C_{0}(\mathbb{R})$. It is independent of the choice of orientation.
Assuming that in terms of local coordinates $\{z_{1},z_{2},...,z_{n}\}$ on $V$ the form $Q$ is given by ${Q}=z_{1}^{2}+...+z_{n-1}^{2}-z_{n}^{2}$, then the form $\omega$ is given on the open set $U=\{p\in C_{0}|z_{n}(p)\neq 0\}$ by
$\omega=\frac{\pm i}{2z_{n}}dz_{1}\wedge...\wedge dz_{n-1}.$
The real cone $C_{0}(\mathbb{R})$ has a decomposition into connected components
$C_{0}(\mathbb{R})=C_{+}(\mathbb{R})\sqcup C_{-}(\mathbb{R}),$
which is determined by the sign of $z_{n}$.
On each of the half cones $C_{+}(\mathbb{R}),C_{-}(\mathbb{R})$ the canonical positive density $|\omega|$ is given by
$|\omega|=\frac{1}{2|z_{n}|}|dz_{1}\wedge...\wedge dz_{n-1}|.$
We can express $z_{n}$ on $C_{\pm}(\mathbb{R})$ as $\pm\sqrt{z_{1}^{2}+...+z_{n-1}^{2}}$.
4.1.5 The canonical Hilbert space $\boldsymbol{H}$
We define the canonical Hilbert space $\boldsymbol{H}$ associated with the PLQS $(\mathcal{V},q,w)$ to be the $L^{2}$ space corresponding to ${C}_{0}(\mathbb{R})$ with the density $|\omega|$, explicitly $\boldsymbol{H}=L^{2}(C_{0}(\mathbb{R}),|\omega|)$. The Hilbert space does not depend on the orientation of $(V,Q)$.
The decomposition $C_{0}(\mathbb{R})=C_{+}(\mathbb{R})\sqcup C_{-}(\mathbb{R})$ induces a decomposition of Hilbert spaces
$\boldsymbol{H}=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}.$ (1)
5 The distinguished self-adjoint action of the algebra $\mathcal{D}(C)$ on $\boldsymbol{H}$
In this section we describe a distinguished ”Schwartz” subspace $\boldsymbol{H}^{\infty}$ of $\boldsymbol{H}$ that carries a self-adjoint action of the algebra $\mathcal{D}(C)$. There are different methods to construct this Schwartz subspace. We first construct the group $G=G(\mathcal{V})$ of quantum symmetries of our Lorentzian quadratic space $(\mathcal{V},q)$. This group is an extension of $O(\mathcal{V})$. We describe a canonical extension of the representation of $O(\mathcal{V})$ on $\boldsymbol{H}$ to a unitary representation $\pi$ of $G$ on $\boldsymbol{H}$. We construct $\boldsymbol{H}^{\infty}$ as the subspace of smooth vectors for $\pi$, which we think of as a suitable Schwartz subspace of $\boldsymbol{H}$.
In Subsection 5.1 we start by considering a general notion of an action of a *-algebra on a Hilbert space and in particular symmetric and self-adjoint actions.
In Subsection 5.2 we introduce the intermediate subspace $C_{c}^{\infty}(C_{0}(\mathbb{R}))$ of smooth compactly supported functions on the cone. This subspace is the domain of a symmetric action of the algebra $\mathcal{D}(C)$.
In Subsection 5.3 we recall that $\boldsymbol{H}$ carries a distinguished unitary irreducible representation $\pi$ of the group $G$. We show that the canonical subspace $\boldsymbol{H}^{\infty}$ consisting of smooth vectors for $\pi$ carries a self-adjoint action of $\mathcal{D}(C)$ which is an extension of the above mentioned symmetric action.
Throughout this section we will assume that we are given an $n$-dimensional PLQS $(\mathcal{V},q,w)$ with $n$ being an even integer such that $n>2$.
5.1 Action of a *-algebra on a Hilbert space
In this section we recall the notion of a partial action of a *-algebra on a Hilbert space and discuss some of its properties. We essentially follow [MR283580] (also see [MR333743, MR3722559]).
5.1.1 A partial action of a *-algebra
Recall that a complex *-algebra AA is an associative unital algebra over $\mathbb{C}$ that is equipped with an involution $A\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow A$ which is anti-linear, anti-homomorphism. Explicitly $\forall a,b\in A$ and $\alpha\in\mathbb{C}$,
$(a^{*})^{*}=a,(a+b)^{*}=a^{*}+b^{*},(ab)^{*}=b^{*}a^{*},1^{*}=1,(\alpha a)^{*}=\overline{\alpha}a^{*}.$
The following definition is a generalization of the theory of unbounded operators on a Hilbert space.
Let $A$ be a complex *-algebra and $\boldsymbol{H}$ a complex Hilbert space.
Definition.
A partial action of $A$ on $\boldsymbol{H}$ is a pair $(\rho,\mathcal{W})$ where $\mathcal{W}$ is a linear subspace of $\boldsymbol{H}$, and $\rho:A\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\mathcal{W})$ is an action of the unital associative algebra $A$ on $\mathcal{W}$.
We call $\boldsymbol{W}$ the domain of the action.
For two partial actions $(\rho,\mathcal{W})$ and $(\rho^{\prime},\mathcal{W}^{\prime})$ of the same *-algebra on the same Hilbert space, we write $\rho\preceq\rho^{\prime}$ whenever $\rho^{\prime}$ is an extension of $\rho$.
5.1.2 A dense action of a *-algebra and its adjoint
A partial action $(\rho,\mathcal{W})$ is called a dense action if the domain $\mathcal{W}$ is dense in $\boldsymbol{H}$. For such a dense action, we define another partial action of $A$ on $\boldsymbol{H}$, the adjoint action $(\rho^{\dagger},\mathcal{W}^{\dagger})$ as follows. The domain $\mathcal{W}^{\dagger}$ consists of all $v\in\boldsymbol{H}$ for which for every $a\in A$, the linear functional on $\mathcal{W}$ that is given by
$x\longmapsto\langle\rho(a)x,v\rangle$
is bounded by $C(v)\|x\|$. The adjoint action $\rho^{\dagger}$ is uniquely defined by
$\langle\rho^{\dagger}(a)v,x\rangle=\langle v,\rho(a^{*})x\rangle$
for every $a\in A$, $v\in\mathcal{W}^{\dagger}$ and $x\in\mathcal{W}$. Obviously $\rho\preceq\rho^{\prime}$ implies that $(\rho^{\prime})^{\dagger}\preceq\rho^{\dagger}$.
5.1.3 Symmetric, self-adjoint and essentially self-adjoint actions
Definition (Compare with Sec. II &\& IV of [MR283580]).
  1. 1.
    A dense action $\rho$ of $A$ on a complex Hilbert space $\boldsymbol{H}$ is symmetric if its adjoint $\rho^{\dagger}$ is an extension of $\rho$, that is, if $\rho\preceq\rho^{\dagger}$. In other words this means that
    $\langle\rho(a)x,v\rangle=\langle x,\rho(a^{*})v\rangle,$
    for $a\in A$, $v,x\in\mathcal{W}$.
  2. 2.
    A symmetric dense action $\rho$ is self-adjoint if it is equal to its adjoint, that is, if $\rho=\rho^{\dagger}$.
  3. 3.
    A symmetric dense action $\rho$ is essentially self-adjoint if $\rho^{\dagger}$ is self-adjoint.
Proposition.
Let $\rho$ be a symmetric dense action of $A$ on a complex Hilbert space $\boldsymbol{H}$. The action $\rho$ is essentially self-adjoint if and only if its adjoint is symmetric. Moreover, in that case for every symmetric extension $\rho^{\prime}$ of $\rho$ we have $(\rho^{\prime})^{\dagger}=\rho^{\dagger}$.
The proof is straightforward.
5.1.4 An important example of a self-adjoint action
A unitary representation of a Lie group $G$ gives rise to a dense self-adjoint action of the universal enveloping algebra $A=\mathcal{U}(\mathfrak{g})$ where $\mathfrak{g}$ is the complexification of the Lie algebra $\operatorname{Lie}(G)$.
Let $(\pi,G,\boldsymbol{H})$ be a unitary representation of a Lie group $G$ on a Hilbert space $\boldsymbol{H}$. Let $\boldsymbol{H}^{\infty}$ be the subspace of smooth vectors in $\boldsymbol{H}$ (recall that a vector is smooth if the corresponding orbit map is smooth). The universal enveloping algebra $\mathcal{U}(\mathfrak{g})$ is a complex *-algebra with respect to the unique involution $\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}(\mathfrak{g})$ that is minus the identity on $\operatorname{Lie}(G)$.
Claim.
The natural action $\pi^{\infty}$ of $\mathcal{U}(\mathfrak{g})$ on $\boldsymbol{H}^{\infty}$ is a dense self-adjoint action.
We will prove this claim in Appendix 8.7. Further information on the underlying smooth space of a representation of a Lie group can be found in [BK2014].
5.2 The canonical symmetric action of $\mathcal{D}(C)$ on $\boldsymbol{H}$
As a first step for constructing the Schwartz subspace $\boldsymbol{H}^{\infty}$, in this subsection we introduce a canonical symmetric action of the algebra $\mathcal{D}(C)$ on $\boldsymbol{H}$ with the domain of compactly supported smooth functions on the cone.
5.2.1 The symmetric action
The algebra $\mathcal{D}(C)$ naturally acts on $C^{\infty}(C_{0}(\mathbb{R}))$, the space of smooth complex-valued functions on $C_{0}(\mathbb{R})$. We denote this action by $\rho^{\infty}$.
The subspace $C_{c}^{\infty}(C_{0}(\mathbb{R}))$ of compactly supported functions is a dense $\rho^{\infty}$-stable subspace of $C^{\infty}(C_{0}(\mathbb{R}))$.
We denote the action of the algebra $\mathcal{D}(C)$ on $C_{c}^{\infty}(C_{0}(\mathbb{R}))$ by $\rho_{c}$.
Claim.
For every $d\in\mathcal{D}(C_{0})$ there is a unique $d^{*}\in\mathcal{D}(C_{0})$ such that
$\langle\rho_{c}(d)f_{1},f_{2}\rangle=\langle f_{1},\rho_{c}(d^{*})f_{2}\rangle,\quad\forall f_{1},f_{2}\in C^{\infty}_{c}(C_{0}(\mathbb{R})).$
Proof.
Differential operators on $C_{0}$ are described by their restrictions to affine open subsets. We will construct a family $\{U_{P}\}$ of such subsets with the following properties
  1. 1.
    Every $U_{P}$ is a dense affine open subset defined over $\mathbb{R}$.
  2. 2.
    The family $\{U_{P}\}$ is closed under intersections and covers $C_{0}$.
  3. 3.
    For every set $U_{P}$, the algebra of differential operators $\mathcal{D}(U_{P})$ satisfies the following condition;
    ()(*) For every $d\in\mathcal{D}(U_{P})$ there is a unique $d^{*}\in\mathcal{D}(U_{P})$ such that
    $\langle\rho_{c}(d)f_{1},f_{2}\rangle=\langle f_{1},\rho_{c}(d^{*})f_{2}\rangle,\quad\forall f_{1},f_{2}\in C^{\infty}_{c}(U_{P}(\mathbb{R})).$
Since $\mathcal{D}(C_{0})$ can be glued out of the algebras $\{\mathcal{D}(U_{P})\}$, a standard gluing argument implies the claim.
Let $P$ be a non-zero polynomial on $C$ defined over $\mathbb{R}$ such that $P(0)=0$. We define the set $U_{P}$ by $U_{P}=C\setminus\{\text{zeroes of }P\}$. Conditions (1.) and (2.) are obviously satisfied. Uniqueness in condition (3.) follows from the fact that $U_{P}(\mathbb{R})$ is Zariski dense in $U_{P}$. Existence of an operator $d^{*}$ follows from the fact that the algebra $\mathcal{D}(U_{P})$ is generated by vector fields given by element in $\operatorname{Lie}(O(q))$ and by regular functions $\mathcal{O}(U_{P})$. For all such generators $d$ the existence of the operator $d^{*}$ is clear. ∎
Remark.
Note that since $\mathcal{D}(C)=\mathcal{D}(C_{0})$ the claim says that there is a unique structure of a * algebra of $\mathcal{D}(C)$ such that the action of $\mathcal{D}(C)$ on $C^{\infty}_{c}(C_{0}(\mathbb{R}))$ is a symmetric dense action $\rho_{c}$ of $\mathcal{D}(C)$ on $\boldsymbol{H}$.
5.2.2 The induced real structure on the Lie algebra $\mathfrak{s}$
In this subsection we explain how the symmetric action of $\mathcal{D}(C)$ induces a real structure on $\mathfrak{s}$ which has a natural reductive dual pair.
The Lie algebra $\tilde{\mathfrak{s}}$ is stable under the involution *. On this complex Lie algebra $\tilde{\mathfrak{s}}$ and its subalgebra $\mathfrak{s}$, the negative of the involution * defines an antiholomorphic involution
$d\longmapsto-d^{*}.$
The fixed-points sub algebra $\mathfrak{s}^{*}(\mathbb{R}):=\{d\in\mathfrak{s}|d^{*}=-d\}$ is a real form of $\mathfrak{s}$ different then the real form $\mathfrak{s}(\mathbb{R})$ introduced in Subsection 2.2.6. We will show that these two real forms are isomorphic, see details in appendix 8.2.2 and appendix 8.2.3.
Among these two real forms, the main one that we use in this paper is $\mathfrak{s}^{*}(\mathbb{R})$.
5.3 The distinguished self-adjoint extension
In this Subsection our goal is to describe a canonical self-adjoint extension $(\rho,\boldsymbol{H}^{\infty})$ of the symmetric action $(\rho_{c},C_{c}^{\infty}(C_{0}(\mathbb{R})))$. Namely, we describe a subspace $\boldsymbol{H}^{\infty}\subset\boldsymbol{H}$ in-between $C_{c}^{\infty}(C_{0}(\mathbb{R}))$ and $C^{\infty}(C_{0}(\mathbb{R}))$, invariant with respect to the action of the algebra $\mathcal{D}(C)$, such that the action $(\rho,\boldsymbol{H}^{\infty})$ is self-adjoint.
There are several descriptions of this space. We will present a construction that realizes $\boldsymbol{H}^{\infty}$ as the space of smooth vectors in a representation of a “large group” $G=G(\mathcal{V})\supset O(\mathcal{V})$ on the Hilbert space $\boldsymbol{H}$.
5.3.1 A canonical extension of the representation of $O(\mathcal{V})$ on $\boldsymbol{H}$
In this section we construct the group $G=G(\mathcal{V})$ of quantum symmetries of $(\mathcal{V},q)$ which is a group extension of $O(\mathcal{V})$. We describe a distinguished representation $\pi$ of $G$ on the canonical Hilbert space $\boldsymbol{H}$.
The extension $G$ of $O(\mathcal{V})$: Let $(\mathbb{R}^{1,1}=\mathbb{R}^{2},q_{1,1})$ be the standard hyperbolic plane with the quadratic form $q_{1,1}(x,y)=xy$. This is a two-dimensional Lorentzian quadratic space. The standard basis $v_{1}:=(1,0)$, $v_{2}:=(0,1)$ consists of isotropic vectors satisfying $B_{1,1}(v_{1},v_{2})=1$,where $B_{1,1}$ is the corresponding symmetric bilinear form.
[…]
We define the Poincaré subgroup $P=P_{v_{1}}$ of $G$ to be the stabilizer of $v_{1}$ in $G$. Clearly
$L\subset P\subset G.$
Our next goal is to show that there exists a canonical extension of the representation of $L$ on $\boldsymbol{H}$ to a representation of $G$.
We will do it in two steps. First we extend it to a representation of $P$, and then to a representation of $G$.
The extension of the representation to $P$: We let $N$ be the unipotent radical of P{P}. It follows that P=LN{P}=LN as a semidirect product with $N$ a normal subgroup isomorphic to $\mathcal{V}^{*}$. Explicitly, an $L$-equivariant isomorphism of abelian groups $f_{v_{2}}:N\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{V}^{*}$ is given by
$\left(f_{v_{2}}(n)\right)(v):=B(nv_{2},v),$
for any $n\in N$ and $v\in\mathcal{V}$, and where $nv_{2}$ stands for the application of the linear transformation $n$ to $v_{2}$.
The representation of $L$ on $\boldsymbol{H}$ extends to a representation $\pi=\pi_{n}$ of $P$, by defining $\pi(n)$ for $n\in N$ to be the operator of multiplication by the function on $C_{0}(\mathbb{R})$ given by $v\longmapsto e^{iB(nv_{2},v)}$.
By Mackey theory this unitary representation of $P$ is irreducible, also see [KOBAYASHI2003551, Prop. 3.3].
The extension of the representation to $G$: The extension of $\pi_{n}=\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H})\right)$ to a representation of $G$ is based on the fact that $G$ has a distinguished representation $\widetilde{\pi}$ of minimal Gelfand-Kirillov dimension, namely the local theta lift of the trivial representation of $SL_{2}(\mathbb{R})$, see [MR1457244]. Further info on the theta correspondence can be found in [MR546602, MR985172, MR986027].
Theorem.
For any $n\in 2\mathbb{N}$ with $n\geq 4$, the representation $\pi=\pi_{n}:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H}\right)$ can be extended uniquely to a unitary irreducible representation of $G$ that is isomorphic to $\widetilde{\pi}_{n}$.
Proof.
Uniqueness follows from irreducibility of $\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H}\right)$ together with Schur’s lemma.
The hard part of the proof is to show the existence of such an extension. This is proven in [KOBAYASHI2003551, Thm. 3]. ∎
Remark.
1. The existence of the extension of $\pi$ from P{P} to $G$ is highly non-trivial. One small evidence for that is that in the action of the element $I\in G$ there is a non-obvious sign-rule. Explicitly, the action is given by
$\pi(-I)=(-1)^{(n-2)/2}.$
2. For $n>4$ the representation $\widetilde{\pi}_{n}$ is the unique minimal representation of $G$, see [KobayashiMano2011].
3. In this paper we shall only use the Schrödinger model of $\widetilde{\pi}$ which is realized on the canonical Hilbert space $\boldsymbol{H}$. Other important realizations of $\widetilde{\pi}$ exist. Several of them are constructed in [KobayashiMano2011]. Also see [MR2020550, MR1108044, MR1172839].
5.3.2 The smooth subspace of $\pi$ as a self-adjoint extension
We let $\pi:G\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}(\boldsymbol{H})$ be the distinguished unitary irreducible representation on the canonical Hilbert space that was discussed above in Subsection 5.3.1, and let $\boldsymbol{H}^{\infty}$ be the canonical subspace of smooth vectors of $\pi$ (the Schwartz space of $\boldsymbol{H}$).
Proposition.
  1. 1.
    $C^{\infty}_{c}(C_{0}(\mathbb{R}))\subseteq\boldsymbol{H}^{\infty}\subseteq C^{\infty}(C_{0}(\mathbb{R}))$.
  2. 2.
    $\boldsymbol{H}^{\infty}$ is invariant under the action $\rho^{\infty}$ of $\mathcal{D}(C)$.
  3. 3.
    The action $\rho=\rho^{\infty}$ of the algebra $\mathcal{D}(C)$ on $\boldsymbol{H}^{\infty}$ is self-adjoint.
Note that in particular the proposition says that the action of the algebra $\mathcal{D}(C)$ on $\boldsymbol{H}^{\infty}$ is a self-adjoint extension of the action on $C^{\infty}_{c}(C_{0}(\mathbb{R}))$.
Proof.
The first part of the proposition is proven in [KobayashiMano2011, Sec. 2.5]. From the explicit formulas by which $\operatorname{Lie}(G)$ acts on $\boldsymbol{H}^{\infty}$ via differential operators on $C_{0}(\mathbb{R})$ e.g. see [KobayashiMano2011, Sec. 2.3], it is evident that the self-adjoint action $\pi^{\infty}:\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{H}^{\infty})$ where $\mathfrak{g}$ is the complexification of the Lie algebra $\operatorname{Lie}(G)$ (as in Subsection 5.1.4 ) factors through a quotient surjective map $j:\mathcal{U}(\mathfrak{g})\twoheadrightarrow\mathcal{D}(C)$. As a result we obtain an action of $\mathcal{D}(C)$ on $\boldsymbol{H}^{\infty}$ and in particular $\boldsymbol{H}^{\infty}$ is invariant under the action $\rho^{\infty}$. We denote the obtained action on $\boldsymbol{H}^{\infty}$ by $\rho$.
Since $j$ is a morphism of *-algebras and $\pi^{\infty}$ is self-adjoint $\rho$ is self-adjoint. ∎
Remark.
1. It is easy to directly calculate $j$ on $\operatorname{Lie}({P})$. In addition, the map $j$ is $(O(Q)\times\mathbb{C}^{\times})$-equivariant. Using Proposition 8.2.1 it follows that there is a unique $(O(Q)\times\mathbb{C}^{\times})$-equivariant extension of $j|_{\operatorname{Lie}({P})}$ to the complexification of $\operatorname{Lie}(G)$. In addition $j(\operatorname{Lie}(G))$ is equal to the real form $\mathfrak{s}^{*}(\mathbb{R})$ introduced in Subsection 5.2.2.
2. The decomposition $\boldsymbol{H}=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}$ is stable under the connected component of the identity $G_{0}\subset G$, see [KobayashiMano2011, p.19]. This implies that $\boldsymbol{H}^{\infty}=\boldsymbol{H}^{\infty}_{+}\oplus\boldsymbol{H}^{\infty}_{-}$, where $\boldsymbol{H}^{\infty}_{\pm}:=\boldsymbol{H}_{\pm}\cap\boldsymbol{H}^{\infty}$ and moreover $\boldsymbol{H}^{\infty}_{\pm}$ are stable under $\rho$.
6 The model
In this section we suggest a new model for the quantum-mechanical system of the hydrogen atom.
6.1 The main ingredients
In physics, a quantum-mechanical system such as the hydrogen atom, is described by a Hilbert space and a distinguished densely defined (unbounded) self-adjoint operator called the Schrödinger operator. The spectrum of the Schrödinger operator and the expansion of a given state in terms of eigenstates of the Schrödinger operator determine the time-evolution of the system via the Schrödinger equation. Other physically relevant (measurable) quantities are given via densely defined (unbounded) symmetric operators (these are called observables).
In the following subsections we list the four ingredients of our suggested model. Our model is associated with an $n$-dimensional PLQS $(\mathcal{V},q,w)$, and a real positive number $\kappa$. We fix such $(\mathcal{V},q,w)$ and $\kappa$ throughout this section.
6.1.1 The underlying Hilbert space
The first ingredient is the canonical Hilbert space associated with the $n$-dimensional PLQS $(\mathcal{V},q,w)$
$\boldsymbol{H}=L^{2}({C}_{0}(\mathbb{R}),|\omega|),$
as in Section 4.1.5.
6.1.2 The algebra of operators
The second ingredient is the *-algebra of differential operators on the cone,
$\mathcal{D}(C).$
6.1.3 The self-adjoint dense action
The third ingredient is the self-adjoint dense action $\rho$ of the *-algebra $\mathcal{D}(C)$ on $\boldsymbol{H}$ with domain $\boldsymbol{H}^{\infty}$.
Note that the algebra $\mathcal{D}(C)$ and its canonical self-adjoint action $(\rho,\boldsymbol{H}^{\infty})$ encode the $G\simeq O(n,2)$ symmetry of the system.
The subspace $\boldsymbol{H}^{\infty}$ replaces the boundary conditions that are typically used in Physics for the quantum-mechanical system of the hydrogen atom.
6.1.4 The Schrödinger family
The fourth ingredient is a one-parameter family in $\mathcal{D}(C)$ called the Schrödinger family.
Recall that with the linear functional $w$ given in our PLQS $(\mathcal{V},q,w)$ we associate a second-order differential operator on the cone $\Psi^{Q}(w)$.
The operators
$e=e_{w}:=iw,\quad h,\quad f=f_{w}:=i\Psi^{Q}(w),$
form an $\mathfrak{sl}_{2}$-triple in $\mathfrak{s}\subset\mathcal{D}(C)$ satisfying the standard $\mathfrak{sl}_{2}$ commutation relations:
$[e,f]=h,\hskip 2.84526pt[h,f]=-2f,\hskip 2.84526pt[h,e]=2e.$
Definition.
Under our assumptions of a fixed PLQS $(\mathcal{V},q,w)$, and a fixed $\kappa\in\mathbb{R}$, for every $\lambda\in\mathbb{C}$, we define a second-order differential operator on the cone
$\boxed{S_{(q,w,\kappa)}(\lambda)=S_{\kappa}(\lambda):=\kappa-i\left(\lambda e+f\right).}$
We call $S_{(q,w,\kappa)}$ the Schrödinger family associated with $(\mathcal{V},q,w,\kappa)$.
Remark.
[…]
Similarly, for any $\mathfrak{sl}_{2}$-triple $\{E,H,F\}$ in any Lie algebra $\mathfrak{g}$ we define a one-parameter family in $\mathcal{U}(\mathfrak{g})$ (the abstract Schrödinger family) by
$S^{Ab}_{\kappa}(\lambda):=\kappa-i\left(\lambda E+F\right),\quad\lambda\in\mathbb{C}.$
6.2 Motivation for the definition of the Schrödinger family
In this paragraph we compare our Schrödinger family as a family of differential operators on the upper half cone with the Schrödinger operator of the hydrogen atom in physics.
In Physics, the Schrödinger operator $H_{phys}$ on $\mathbb{R}^{3}\setminus\{0\}$ is given by
$H_{phys}=-\Delta-\frac{\kappa}{r}$}
where $\Delta$ is the Laplacian, $r$ the radius function and $\kappa$ is a positive constants. The corresponding time-independent Schrödinger equation on $\mathbb{R}^{3}\setminus\{0\}$ is given by
$H_{phys}\psi=E\psi,$}
where $E$ is the energy (the eigenvalue of $H_{phys}$).
Comparing the above with our suggested model we can choose the $4$-dimensional PLQS in which $\mathcal{V}=\mathbb{R}^{4}$, $q(x,y,z,w)=x^{2}+y^{2}+z^{2}-w^{2}$, and the fixed linear functional in our PLQS is the linear functional $w$ (determined by the coordinate $w$).
Projection from the upper half cone
${C}_{+}(\mathbb{R})=\{(x,y,z,w)\in C_{0}(\mathbb{R})|w>0\}$
onto $\mathbb{R}^{3}\setminus\{0\}$ is a diffeomorphism that enables us to express the Schrödinger family $S_{(q,w,\kappa)}(\lambda=E)$ in terms of the Cartesian coordinates $x,y,z$, as
$r\Delta+\kappa+\lambda r=r\left(-H_{phys}+E\right).$
That is, on the space of smooth functions on $\mathbb{R}^{3}\setminus\{0\}$, the operator $S_{(q,w,\kappa)}(E)$ is proportional to $(H_{phys}-E)$ and both have the same kernel.
7 Spectrum of the Schrödinger family
In this section we discuss the spectrum of a family of operators on a topological vector space.
We compare the spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}$ with the spectrum of the Schrödinger operator of the Hydrogen atom in physics. We show that the later coincides with the spectrum of the Schrödinger family in $\boldsymbol{H}_{+}^{\infty}$ the subspace consisting of functions that are supported on the ”positive half cone”.
7.1 Invertibility and spectrum
In this subsection we define the spectrum of a family of continuous operators on topological vector space.
Our goal is to calculate the spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}$. In this setting the Schrödinger family acts as a family of operators within a representation of $SL_{2}(\mathbb{R})$ that commute with $SO(n)$. We show that it is enough to calculate the spectrum within each $SO(n)$-isotypic component.
7.1.1 The spectrum of a family of operators
In functional analysis, given a continuous operator $T$ on a dense linear subspace $\boldsymbol{W}$ of a topological vector space $\boldsymbol{H}$ (typically $\boldsymbol{H}$ is a Hilbert space) the spectrum of $T$ is defined as the collection of all $\lambda\in\mathbb{C}$ for which the operator $T-\lambda:\boldsymbol{W}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}$ does not have a continuous inverse, see e.g., Section VI3 of [reed]. In physics, given an operator $T$ as above, one is usually interested in a spectral decomposition of $\boldsymbol{H}$ with respect to $T$. Sometimes there are not enough eigenvectors in $\boldsymbol{H}$ for spectral decomposition and a natural extension is the dual space $\boldsymbol{W}^{*}$. This motivates the following definition.
Let $\boldsymbol{W}$ be a topological vector space. Consider a family $\mathcal{F}$ of continuous operators on $\boldsymbol{W}$ parametrized by a set $\Lambda$ (i.e., a map $\mathcal{F}:\Lambda\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{W})$).
Definition.
Let $\mathcal{F}:\Lambda\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{W})$ be a family of continuous linear operators. For every $\lambda\in\Lambda$ we define the $\lambda$-solutions space (of $\mathcal{F}$) in $\boldsymbol{W}$, to be the kernel of $\mathcal{F}(\lambda)^{*}$, where $\mathcal{F}(\lambda)^{*}:\boldsymbol{W}^{*}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{W}^{*}$ is the adjoint of $\mathcal{F}(\lambda)$. We denote this space by $\operatorname{Ker}(\mathcal{F}(\lambda);\boldsymbol{W}^{*})$.
We define $\operatorname{Spec}_{\boldsymbol{W}}\{\mathcal{F}\},$ the spectrum of the family $\mathcal{F}$, to be the collection of all $\lambda\in\Lambda$ with nonzero solutions space. That is,
$\operatorname{Spec}_{\boldsymbol{W}}\{\mathcal{F}\}=\{\lambda\in\Lambda|\operatorname{Ker}(\mathcal{F}(\lambda);\boldsymbol{W}^{*})\neq 0\}.$
Example.
For a given continuous operator $T$ on a topological vector space $\boldsymbol{W}$, we obtain a family $\mathcal{F}:\mathbb{C}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{W})$ via $\mathcal{F}(\lambda)=T-\lambda$. Then $\operatorname{Spec}_{\boldsymbol{W}}\{\mathcal{F}\}$ is equal to the point-spectrum of the adjoint $T^{*}$, and $\operatorname{Ker}(\mathcal{F}(\lambda);\boldsymbol{W}^{*})$ is equal to the corresponding eigenspace.
When $\boldsymbol{W}=C^{\infty}_{c}(\mathbb{R})$, the space of smooth compactly supported functions on the line equipped with the standard topology and $T=\frac{d}{dx}$, $\operatorname{Spec}_{\boldsymbol{W}}\{\mathcal{F}\}=\mathbb{C}$ and $\operatorname{Ker}(\mathcal{F}(\lambda);\boldsymbol{W}^{*})$ consists of the distributional solutions of the ODE $\frac{d}{dx}\varphi=\lambda\varphi$. Explicitly, $\operatorname{Ker}(\mathcal{F}(\lambda);\boldsymbol{W}^{*})$ is spanned by the distribution $f\longmapsto\int_{\mathbb{R}}e^{\lambda x}f(x)dx$.
Finally, when $\boldsymbol{W}$ is the space of Schwartz class functions on $\mathbb{R}$ and $T=\frac{d}{dx}$, the spectrum reduces to $i\mathbb{R}$.
7.1.2 The spectrum of the Schrödinger family within a representation of $SL_{2}(\mathbb{R})$
For any smooth representation $(\pi,\boldsymbol{W})$ of $SL_{2}(\mathbb{R})$, $\pi(S^{Ab}_{\kappa})$ is a family of operators in $\operatorname{End}(\boldsymbol{W})$ parameterized by $\mathbb{C}$.
The following lemma will be needed later on. We omit its simple proof.
Lemma.
Let $(\pi,\boldsymbol{W})$ be a smooth representation of $SL_{2}(\mathbb{R})$ and assume that $\kappa>0$. Then
$\operatorname{Spec}_{\boldsymbol{W}}\{\pi(S^{Ab}_{\kappa})\}=\kappa^{2}\left(\operatorname{Spec}_{\boldsymbol{W}}\{\pi(S^{Ab}_{1})\}\right).$
7.1.3 The spectrum of the Schrödinger family within a representation of $SL_{2}(\mathbb{R})\times SO(n)$
We will be interested in the spectrum of the Schrödinger family within a representation of $SL_{2}(\mathbb{R})\times SO(n)$, mainly in $\boldsymbol{H}^{\infty}$. In this section we explain that in this setting calculating the spectrum and solutions within each $SO(n)$-isotypic component is enough to determine the spectrum and solutions in the ambient topological vector space.
Given a continuous representation $\pi:SL_{2}(\mathbb{R})\times SO(n)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{Aut}(\boldsymbol{W})$ of $SL_{2}(\mathbb{R})\times SO(n)$ on a topological vector space $\boldsymbol{W}$, and an irreducible (unitary) representation $\tau:SO(n)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{Aut}(\boldsymbol{X}_{\tau})$ of $SO(n)$ on a finite-dimensional Hilbert space $\boldsymbol{X}_{\tau}$ we denote the canonical $SO(n)$-isotypic subspace of $\boldsymbol{W}$ corresponding to $\tau$ by $\boldsymbol{W}_{\tau}$. Each $\boldsymbol{W}_{\tau}$ carries a representation of $SL_{2}(\mathbb{R})$.
Claim.
  1. 1.
    For any continuous representation $\pi:SL_{2}(\mathbb{R})\times SO(n)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{Aut}(\boldsymbol{W})$, any $\lambda\in\mathbb{C}$, and any irreducible representation $\tau:SO(n)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{Aut}(\boldsymbol{X}_{\tau})$, the $\tau$-isotypic component of the $\lambda$-solution space in $\boldsymbol{W}$ is isomorphic to the $\lambda$ solution space in $\boldsymbol{W}_{\tau^{*}}$. That is
    $\left(\operatorname{Ker}(S^{Ab}_{\kappa}(\lambda);\boldsymbol{W}^{*})\right)_{\tau}=\operatorname{Ker}(S^{Ab}_{\kappa}(\lambda);\boldsymbol{W}_{\tau^{*}}^{*}),$
    where $\tau^{*}$ is the contragradient of $\tau$.
  2. 2.
    The union of spectra for the Schrödinger family on all of the $SO(n)$-isotypic peices is equal to the spectrum of the Schrödinger family in $\boldsymbol{W}$. That is,
    $\operatorname{Spec}_{\boldsymbol{W}}\{S^{Ab}_{\kappa}\}=\bigcup_{\tau\in\widehat{SO(n)}}\operatorname{Spec}_{\boldsymbol{W}_{\tau}}\{S^{Ab}_{\kappa}\},$
    where $\widehat{SO(n)}$ stands for the space of equivalence classes of irreducible representations of $SO(n)$.
Verification of this claim is straightforward.
Remark.
We note that $\bigoplus_{\tau\in\widehat{SO(n)}}\boldsymbol{W}_{\tau}$ is dense in $\boldsymbol{W}$.
7.2 The spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}$
In this section we calculate the spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}$. We show that the spectrum of the Schrödinger family in $\boldsymbol{H}^{\infty}_{+}$ coincides with the spectrum that is calculated in Physics in the space of functions (or distributions) on $L^{2}(\mathbb{R}^{3})$ satisfying certain boundary conditions.
We fix $n=4$ to be in agreement with the physical system of the hydrogen atom. The presented results hold with appropriate adjustments for any $n\in 2\mathbb{N}$ such that $n\geq 4$.
7.2.1 The main result
Theorem.
Let $(\mathcal{V},q,w)$ be a $4$-dimensional PLQS, let $\kappa$ be a fixed positive number, and let $\boldsymbol{H}$ be the corresponding canonical Hilbert space.
$\operatorname{Spec}_{\boldsymbol{H}_{+}^{\infty}}\{S^{Ab}_{\kappa}\}=\left\{-\frac{\kappa^{2}}{4n^{2}}|n\in\mathbb{N}\right\}\cup[0,\infty)$
$\operatorname{Spec}_{\boldsymbol{H}_{-}^{\infty}}\{S^{Ab}_{\kappa}\}=(0,\infty).$
The representation $(\pi,G(\mathcal{V}),\boldsymbol{H})$ gives rise to a representation of $SL_{2}(\mathbb{R})\times SO(3)$ on $\boldsymbol{H}$. We shall prove the above theorem using the representation $(\pi,SL_{2}(\mathbb{R})\times SO(3),\boldsymbol{H})$ and Claim 7.1.3 by reducing the spectrum calculations to calculations within each $SO(3)$-isotypic component. This is stated precisely in Proposition 7.2.6 below. Each $SO(3)$-isotypic component carries a discrete series representation of $SL_{2}(\mathbb{R})$. We explain how the above mentioned fact together with Casselman-Wallach theorem reduces the problem of calculating the spectrum of the Scrödinger family in $\boldsymbol{H}^{\infty}$ to a problem that can be handled completely withing representation theory of $SL_{2}(\mathbb{R})$.
7.2.2 The $(SL_{2}(\mathbb{R})\times SO(3))$ representation on $\boldsymbol{H}$
Let $(\pi,G(\mathcal{V}),\boldsymbol{H})$ be the unitary irreducible representation associated with the given PLQS $(\mathcal{V},q,w)$ that was discussed in Subsection 5.3.1.
To simplify our discussion we will identify $\mathcal{V}$ with $\mathbb{R}^{4}$, the Lorentzian quadratic form $q$ with the standard Lorentzian quadratic form $q(x,y,z,w)=x^{2}+y^{2}+z^{2}-w^{2}$, and the fixed linear functional $w$ with the coordinate $w$.
We let ${K}_{w}=SO(3)$ be the connected component of the identity in the stabilizer of $w$ in $O(q)=O(3,1)$. The connected component of the identity in the centralizer of $SO(3)$ in $G(\mathcal{V})$ is isomorphic to $PSL_{2}(\mathbb{R})$. Hence there is a unique homomorphism $\xi:SL_{2}(\mathbb{R})\times SO(3)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow G(\mathcal{V})$ such that;
  1. 1.
    Restricted to $SO(3)$, $\xi$ is the identity.
  2. 2.
    For the elements
    $E:=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},H:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
    in $\mathfrak{sl}_{2}(\mathbb{R})$, as differential operators on the cone
    $\pi(\xi(E))=iw,\quad\pi(\xi(H))=h=2\chi+2.$
Remark.
It can be shown that $\pi(\xi(\mathfrak{sl}_{2}(\mathbb{R})))=\mathfrak{l}_{w}\cap\mathfrak{s}^{*}(\mathbb{R})$, $\pi(\xi(\mathfrak{so}(3)))=\mathfrak{k}_{w}\cap\mathfrak{s}^{*}(\mathbb{R}))$, where $\mathfrak{l}_{w}$ and $\mathfrak{k}_{w}$ are as in Section 3.0.2. Explicitly,
$\pi(\xi(\mathfrak{so}(3)))=\operatorname{Span}_{\mathbb{R}}\{x\partial_{y}-y\partial_{x},y\partial_{z}-z\partial_{y},z\partial_{x}-x\partial_{z}\}$
$\pi(\xi(\mathfrak{sl}_{2}(\mathbb{R})))=\operatorname{Span}_{\mathbb{R}}\{e:=iw,f:=i\Psi^{Q}(w),h=2\chi+2\}.$
Moreover, Proposition 8.2.3 and Proposition 3.0.2 imply that the pair $(\pi(\xi(\mathfrak{sl}_{2}(\mathbb{R}))),\pi(\xi(\mathfrak{so}(3))))$ is a dual pair in $\mathfrak{s}^{*}(\mathbb{R})$.
7.2.3 The $(SL_{2}(\mathbb{R})\times SO(3))$-isotypic decomposition of $\boldsymbol{H}_{\pm}$
For $\ell\in\mathbb{Z}_{\geq 0}$ we denote by $\mathcal{Y}^{\ell}$ the subspace of $L^{2}(S^{2})$ consisting of Spherical Harmonics of degree $\ell$. Each $\mathcal{Y}^{\ell}$ is an irreducible representation of $SO(3)$, and up to equivalence there are no others.
The $SO(3)$-isotypic component of $\boldsymbol{H}$ corresponding to $\mathcal{Y}^{\ell}$ will be denoted by $\boldsymbol{H}_{\ell}$ and of $\boldsymbol{H}^{\infty}$ by $\boldsymbol{H}^{\infty}_{\ell}$.
We denote by $D^{\pm}_{2\ell+2}$ the irreducible unitary discrete series representation of $SL_{2}(\mathbb{R})$ with $SO(2)$-types that are given by $\pm\{2\ell+2,2\ell+4,$...\}, respectively. More concretely, these integers are the eigenvalues of the operator $H_{c}:=i\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$, the generator for the compact one-parameter subgroup $SO(2)$ of $SL_{2}(\mathbb{R})$.
We set $D_{2\ell+2}:=D^{+}_{2\ell+2}\oplus D^{-}_{2\ell+2}$.
The following claim about the action of $(SL_{2}(\mathbb{R})\times SO(3))$ on $\boldsymbol{H}^{\infty}$ can be verified directly.
Claim.
On $\boldsymbol{H}^{\infty}$ the actions of the Casimirs of $\mathfrak{sl}_{2}(\mathbb{R})$ and $\mathfrak{s0}(3)$ coincide.
The claim above together with the known classification of the unitary dual of $SL_{2}(\mathbb{R})$ imply that for $\ell>0$ the only representation of $SL_{2}(\mathbb{R})$ that can appear in $\boldsymbol{H}_{\ell}$ are $D^{\pm}_{2\ell+2}$ and those that can appear in $\boldsymbol{H}_{0}$ are $D^{\pm}_{2}$ and the trivial representation. The next proposition state explicitly which representations of $SL_{2}(\mathbb{R})\times SO(3)$ appear in $\boldsymbol{H}_{\ell}$.
Proposition.
As representations of $SL_{2}(\mathbb{R})\times SO(3)$ on Hilbert spaces, for every $\ell\in\mathbb{N}_{0}$,
$\boldsymbol{H}_{\ell}\cong D_{2\ell+2}\otimes_{\mathbb{C}}\mathcal{Y}^{\ell},$ (2)
$(\boldsymbol{H}_{\pm})_{\ell}\cong D^{\pm}_{2\ell+2}\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}$
This follows from Sections 1.3 and 3.1 of [MR2401813].
Remark.
The multiplicities of $D^{\pm}_{2\ell+2}\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}$ in the canonical Hilbert space $\boldsymbol{H}$ can also be deduced by calculations similar to those that are given below in the proof of Proposition 7.2.6 in the case of $\lambda=0$.
7.2.4 Generalized spherical coordinates on the cone
The cone ${C}_{0}(\mathbb{R})$ is a double cover of $\mathbb{R}^{3}\setminus\{0\}$, this motivates us to use new better suited coordinates on ${C}_{0}(\mathbb{R})$.
We introduce generalized spherical coordinates on the cone via
$\mathbb{R}\setminus\{0\}\times S^{2}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow{C}_{0}(\mathbb{R})$
$(t,u)\longmapsto(tu,t).$
This leads to an isomorphism
$\boldsymbol{H}\cong L^{2}\left(\mathbb{R}\setminus\{0\},\frac{1}{2}|t|dt\right)\otimes_{\mathbb{C}}L^{2}(S^{2}),$
respecting the standard actions of $SO(3)$. From this we see that for any $\ell\in\mathbb{N}_{0}$,
$\boldsymbol{H}_{\ell}\cong L^{2}\left(\mathbb{R}\setminus\{0\},\frac{1}{2}|t|dt\right)\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}.$
Proposition.
The action of $\mathfrak{sl}_{2}(\mathbb{R})$ on $\boldsymbol{H}^{\infty}_{\ell}$ in terms of generalized spherical coordinates is given by
$\pi_{\ell}(E)=i{t},$
$\pi_{\ell}(H)=2t\partial_{t}+2,$
$\pi_{\ell}(F)=i\left(t\partial_{t}^{2}+2\partial_{t}-\frac{\ell(\ell+1)}{t}\right).$
The proof follows from direct calculations using claim 7.2.3.
7.2.5 Implications of the Casselman-Wallach theorem
In this paragraph we explain how Casselman-Wallach globalization theorem allows us to do all the relevant calculations for determining the spectrum of the Schrödinger family completely within representation theory of $SL_{2}(\mathbb{R})$. However, in this paper we solve this problem working with these representations of $SL_{2}(\mathbb{R})$ as they concretely appear inside the Hilbert space $\boldsymbol{H}$ of functions on the cone.
By Casselman-Wallach globalization theorem [Wallach3, Casselman, Wallach2], any two admissible smooth moderate growth Fréchet representations of the same reductive Lie group that are infinitesimally equivalent must be isomorphic as topological representations.
The topological vector space $\boldsymbol{H}^{\infty}$ is admissible smooth moderate growth Fréchet representation of $G\simeq O(4,2)$. The subspaces $(\boldsymbol{H}_{\pm}^{\infty})_{\ell}$ are automatically smooth moderate growth Fréchet representation of $SL_{2}(\mathbb{R})\times SO(3)$. By Proposition 7.2.3 these representations are also admissible, hence
$(\boldsymbol{H}^{\infty}_{+})_{\ell}\cong(D^{+}_{(2\ell+2)})^{\infty}\otimes_{\mathbb{C}}\mathcal{Y}^{\ell},\quad(\boldsymbol{H}^{\infty}_{-})_{\ell}\cong(D^{-}_{-(2\ell+2)})^{\infty}\otimes_{\mathbb{C}}\mathcal{Y}^{\ell},$
where on the right hand side, smoothness is with respect to the action of $SL_{2}(\mathbb{R})$. In appendix 8.5 we explicitly describe the isotypic components $(\boldsymbol{H}^{\infty}_{\pm})_{\ell}$.
7.2.6 The spectrum of the Schrödinger family in $\boldsymbol{H}_{\pm}^{\infty}$
Proposition.
Let $(\mathcal{V},q,w)$ be a $4$-dimensional PLQS, let $\kappa$ be a fixed positive number, and let $\boldsymbol{H}$ be the corresponding canonical Hilbert space. Fix $\ell\in\mathbb{Z}_{\geq 0}$.
  1. 1.
    $\operatorname{Spec}_{(\boldsymbol{H}_{+}^{\infty})_{\ell}}\{S_{\kappa}\}=\left\{-\frac{\kappa^{2}}{4n^{2}}\middle|\ell<n\in\mathbb{N}\right\}\cup[0,\infty).$
    $\operatorname{Spec}_{(\boldsymbol{H}_{-}^{\infty})_{\ell}}\{S_{\kappa}\}=(0,\infty).$
  2. 2.
    For $\lambda\in\mathbb{C}$, the kernel of the operator $S_{\kappa}(\lambda)$ in $((\boldsymbol{H}_{+}^{\infty})_{\ell})^{*}$ is nonzero if and only if one of the following hold:
    1. (a)
      $0\leq\lambda$.
    2. (b)
      $\lambda=-\frac{\kappa^{2}}{4n^{2}}$ for some $n\in\mathbb{N}$ such that $\ell<n$.
    Moreover, in these cases as a representation of $SO(3)$ the kernel is isomorphic to $\mathcal{Y}^{\ell}$.
  3. 3.
    For $\lambda\in\mathbb{C}$, the kernel of the operator $S_{\kappa}(\lambda)$ in $((\boldsymbol{H}_{-}^{\infty})_{\ell})^{*}$ is nonzero if and only if $\lambda>0$. Moreover, in these cases as a representation of $SO(3)$ the kernel is isomorphic to $\mathcal{Y}^{\ell}$.
The above proposition and Claim 7.1.3 prove Theorem Theorem.
Proof.
In Proposition 8.6 in the appendix we show that in any admissible smooth Fréchet representation of $SL_{2}(\mathbb{R})$ the spectrum of the Schrödinger family is always contained in $\mathbb{R}$. In particular this holds for $(\boldsymbol{H}_{\pm}^{\infty})_{\ell}$. Taking this into account and using Lemma 7.1.2, it is enough to determine for which $\lambda\in\mathbb{R}$ the kernel of the operator $S:=S_{1}(\lambda)$ in the space $\left((\boldsymbol{H}_{\pm}^{\infty})_{\ell}\right)^{*}$ is nonzero, and in these cases to calculate its dimension.
We fix $\lambda\in\mathbb{R}$. Our strategy for determining the kernel of $S$ is as follows. We find another, more simple, operator $T$ that is conjugate to $S$ by the action of some element $A\in SL_{2}(\mathbb{R})$.
Then $A$ defines an isomorphism $\operatorname{Ker}(T)\simeq\operatorname{Ker}(S)$, so it is enough to compute the dimension of $\operatorname{Ker}(T)$.
We treat the three cases of negative $\lambda$, $\lambda=0$, and positive $\lambda$, separately.
The case of $\lambda<0$. We set $\nu:=\sqrt{|\lambda|}$ a positive square root. In this case there exists an element $A\in SL_{2}(\mathbb{R})$ such that $\operatorname{Ad}_{A}(S)=1+i\nu E-i\nu F$. Set $T=1+i\nu E-i\nu F$. As a result the kernel of the operator $S$ in $\left((\boldsymbol{H}_{\pm}^{\infty})_{\ell}\right)^{*}$ is isomorphic to the kernel of the operator $T$ in the same space.
This last kernel is equal to the eigenspace of the operator i(FE)i(F-E) in $\left((\boldsymbol{H}_{\pm}^{\infty})_{\ell}\right)^{*}$ corresponding to the eigenvalue $\nu^{-1}$. The one-parameter subgroup generated by FEF-E inside $SL_{2}(\mathbb{R})$ is $SO(2)$. The eigenvalues of FEF-E in any representation of $SL_{2}(\mathbb{R})$ belong to $i\mathbb{Z}$. Moreover the eigenvectors are automatically smooth vectors hence we can solve for the kernel in $(\boldsymbol{H}^{\infty}_{\pm})_{\ell}$ instead of in $\left((\boldsymbol{H}_{\pm}^{\infty})_{\ell}\right)^{*}$.
It follows from Proposition 7.2.3, that the eigenvalues of i(FE)i(F-E) in $(\boldsymbol{H}^{\infty}_{\pm})_{\ell}$ are $\pm 2(k+\ell+1)$ with $k\in\mathbb{N}_{0}$, and that each eigenspace is isomorphic as a representation of $SO(3)$ to the corresponding irreducible space of spherical harmonics.
Hence for any $\lambda<0$,
  1. 1.
    The kernel of $S$ in $((\boldsymbol{H}_{-}^{\infty})^{*})_{\ell}$ is zero for any $\ell\in\mathbb{Z}_{\geq 0}$.
  2. 2.
    The kernel of $S$ in $((\boldsymbol{H}_{+}^{\infty})^{*})_{\ell}$ is nonzero if and only if $\lambda=-\frac{1}{4n^{2}}$ for some $n\in\mathbb{N}$ such that $\ell<n$. Moreover, for $\lambda=-\frac{1}{4n^{2}}$ with $\ell<n$, the kernel is isomorphic to $\mathcal{Y}^{\ell}$.
The case of $\lambda=0$. In this case there exists an element $A\in SL_{2}(\mathbb{R})$ such that $\operatorname{Ad}_{A}(S)=1+iE$. Set $T=1+iE$. As a result the kernel of the operator $S$ in $\left((\boldsymbol{H}_{\pm}^{\infty})_{\ell}\right)^{*}$ is isomorphic to the kernel of the operator $T$ in the same space. This last kernel is a subspace of the kernel of $T$ in $\mathcal{D}^{\prime}(C_{\pm}(\mathbb{R}))_{\ell}$, where $\mathcal{D}^{\prime}(C_{\pm}(\mathbb{R}))$ stands for the space of distributions on $C_{\pm}(\mathbb{R})$.
In terms of the above-mentioned generalized spherical coordinates the operator $T$ is given by
$\left(\pi_{\ell,\pm}(T)\right)^{*}(\varphi)=(1-t)\varphi,\quad\forall\varphi\in\mathcal{D}^{\prime}(C_{\pm}(\mathbb{R}))_{\ell}.$
Hence its kernel on $\mathcal{D}^{\prime}(C_{+}(\mathbb{R}))_{\ell}$ is a product of the distribution $\delta(t-1)$ by any spherical harmonics from $\mathcal{Y}^{\ell}$, where $\mathcal{Y}^{\ell}$ considered as a space of regular distributions on $S^{2}$. This subspace is already in $(\boldsymbol{H}_{+}^{\infty})^{*}$ and hence isomorphic to the kernel of $S$ in $\left((\boldsymbol{H}_{\pm}^{\infty})_{\ell}\right)^{*}$
On the other hand the kernel of $T$ in $\mathcal{D}^{\prime}(C_{-}(\mathbb{R}))_{\ell}$ is trivial and hence so is the kernel of $S$.
The case of $\lambda>0$. We set $\nu:=\sqrt{\lambda}$ a positive square root. In this case there exists an element $A\in SL_{2}(\mathbb{R})$ such that $\operatorname{Ad}_{A}(S)=1+i\nu H$. Set $T=1+i\nu H$. As a result the kernel of $S$ in $((\boldsymbol{H}_{\pm}^{\infty})^{*})_{\ell}$ is isomorphic to a subspace of the kernel of $T$ in $\mathcal{D}^{\prime}(C_{\pm}(\mathbb{R}))_{\ell}$.
In terms of the above-mentioned generalized spherical coordinates $T$ acts via
$\left(\pi_{\ell,\pm}(T)\right)^{*}(\varphi)=\left(1+2i\nu+2i\nu t\partial_{t}\right)\varphi,\quad\forall\varphi\in\mathcal{D}^{\prime}(C_{\pm}(\mathbb{R}))_{\ell}.$
Hence the solution space in $\mathcal{D}^{\prime}(C_{\pm}(\mathbb{R}))_{\ell}$ consists of regular distributions of the form $\varphi(t,u)=t^{-1+\frac{i}{2\nu}}Y^{\ell}(u)$ for any $Y^{\ell}\in\mathcal{Y}^{\ell}$. Explicitly on $\mathcal{D}^{\prime}(C_{+}(\mathbb{R}))_{\ell}$ using generalized spherical coordinates and the measure on the cone this distribution is given by
$f\in C^{\infty}_{c}((0,\infty)\times S^{2})\longmapsto\langle\varphi,f\rangle:=\int_{0}^{\infty}\int_{S^{2}}f(t,u)t^{-1+\frac{i}{2\nu}}Y^{\ell}(u)tdtdu$
and on $\mathcal{D}^{\prime}(C_{-}(\mathbb{R}))_{\ell}$ by
$f\in C^{\infty}_{c}((-\infty,0)\times S^{2})\longmapsto\langle\varphi,f\rangle:=\int^{0}_{-\infty}\int_{S^{2}}f(t,u)t^{-1+\frac{i}{2\nu}}Y^{\ell}(u)|t|dtdu$
The explicit description of the $SO(3)$-isotypic pieces of $\boldsymbol{H}^{\infty}_{\pm}$ given in Section 8.5 implies that the above mentioned distributions in the kernel of $\left(\pi_{\ell,\pm}(T)\right)^{*}$ can be extended to linear functionals on $(\boldsymbol{H}_{\pm}^{\infty})_{\ell}$. ∎
8 Appendix
8.1 The commutation relations in $\mathfrak{s}$
Recall that the Lie algebra $\tilde{\mathfrak{s}}$ that was introduced in Section 2.2.2 is the direct sum of
$\mathcal{D}^{0,1}(C)\cong\mathcal{D}^{0,1}(V)=\{\tau(v)|v\in V\},$
$\mathcal{D}^{1,0}(C)\cong\mathcal{D}^{1,0}(V)_{I}=\operatorname{span}_{\mathbb{C}}\{L_{u,v}|u,v\in V\}\oplus\mathbb{C}h\oplus\mathbb{C},$
$\mathcal{D}^{2,-1}(C)\cong\mathcal{D}^{2,-1}(V)_{I}=\{\psi^{Q}(\tau(v))|v\in V\}.$
The subalgebras $\mathcal{D}^{0,1}(V)$ and $\mathcal{D}^{2,-1}(V)_{I}$ are commutative and the element $1$ is central. By direct calculations, for any $u,v,w,z\in V$,
$[L_{u,v},L_{w,z}]=B(v,w)L_{uz}-B(u,w)L_{vz}+B(u,z)L_{vw}-B(v,z)L_{uw}$
$[L_{u,v},\tau(w)]=B(v,w)\tau(u)-B(u,w)\tau(v)$
$[L_{u,v},\psi^{Q}(\tau(w))]=B(v,w)\psi^{Q}(\tau(u))-B(u,w)\psi^{Q}(\tau(v))$
$[\psi^{Q}(\tau(v)),\tau(u)]=2L_{v,u}-B(v,u)h$
$[h,\psi^{Q}(\tau(v))]=-2\psi^{Q}(\tau(v))$
$[h,\tau(u)]=2\tau(u)$
$[h,L_{u,v}]=0.$
8.2 The Lie algebra $\mathfrak{s}$ as an orthogonal Lie algebra
8.2.1 The Lie algebra $\mathfrak{s}$ as an orthogonal Lie algebra
We let $(V_{2},Q_{2})$ be the complexification of the standard hyperbolic plane $(\mathbb{R}^{1,1}=\mathbb{R}^{2},q_{1,1})$ from Subsection 5.3.1 with its basis $\{v_{1},v_{2}\}$ of isotropic vectors satisfying $B_{1,1}(v_{1},v_{2})=1$.
[…]
We denote the Lie algebras of $G$, $P$, $L$ and $A$ by $\mathfrak{g}$, $\mathfrak{n}$, $\mathfrak{l}$ and $\mathfrak{a}$ respectively.
In this section we construct a canonical extension of the natural morphism of Lie algebras $\mathfrak{l}=\mathfrak{o}(Q)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ to a morphism of Lie algebras $\psi:\mathfrak{g}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$. This will give us an epimorphism of associative algebras $\psi:\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$.
We construct the embedding in two steps. First we construct a morphism $\mathfrak{p}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ and then we extend it to $\mathfrak{g}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$.
The action of $A$ on $v_{1}\in V_{2}$ defines an isomorphism $\rho:A\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathbb{C}^{\times}$ given by
$av_{1}=\rho(a)v_{1},\quad\forall a\in A.$
The group $G$ acts by conjugation on its Lie algebra $\mathfrak{g}$, and using the homomorphism
$\mathbb{C}^{\times}\xrightarrow{\rho^{-1}}A\lhook\mathrel{\mkern-3.0mu}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow G,$
$\mathbb{C}^{\times}$ also acts on $\mathfrak{g}$.
There is an obvious inclusion of Lie algebras $\iota_{\mathfrak{o}}:\mathfrak{l}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathfrak{g}$. There is also a natural embedding of Lie algebras $\iota_{\mathcal{D}}:\mathfrak{l}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}$, obtained by differentiating the action of $L$ on regular functions on $V$.
We have a natural isomorphism of abelian Lie algebras
$f_{v_{2}}:\mathfrak{n}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow V^{*}\subset\mathcal{D}(C)$
given by
$\left(f_{v_{2}}(n)\right)(v):=B(nv_{2},v),$
for any $n\in N$ and $v\in\mathcal{V}$, and where $nv_{2}$ stands for the application of the linear transformation $n$ to $v_{2}$.
Hence we obtain a unique linear map $\widetilde{\psi}:\mathfrak{p}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ satisfying $\widetilde{\psi}\circ\iota_{\mathfrak{o}}=\iota_{\mathcal{D}}$ and $\widetilde{\psi}|_{\mathfrak{n}}=f_{v_{2}}$. It easily verified that $\widetilde{\psi}$ is an $(L\times\mathbb{C}^{\times})$-equivariant embedding of complex Lie algebras.
Proposition.
There exists a unique $\mathbb{C}^{\times}$-equivariant morphism of complex Lie algebras $\psi:\mathfrak{g}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ extending $\widetilde{\psi}$. In addition $\psi$ is an embedding, its image is equal to $\mathfrak{s}$, and the induced map $\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}(C)$ is an epimorphism.
The proof is left to the reader.
8.2.2 The induced real structure on $\mathfrak{s}$
The construction of $\tilde{\mathfrak{s}}$ out of a complex quadratic space $(V,Q)$ can be repeated over the field of real numbers. Hence, given a real form $(\mathcal{V},q)$ of $(V,Q)$, not necessarily Lorentzian, there is a canonical real form $\mathfrak{s}(\mathbb{R})$ of $\mathfrak{s}$ that is compatible with $(\mathcal{V},q)$ and respects the decomposition $\tilde{\mathfrak{s}}=\mathcal{D}^{0,1}(C)\oplus\mathcal{D}^{1,0}(C)\oplus\mathcal{D}^{2,-1}(C)$. Explicitly, it is given by the direct sum of
$\mathcal{D}^{0,1}(V)(\mathbb{R})=\{\tau^{Q}(v)|v\in\mathcal{V}\},$
$\mathcal{D}^{1,0}(V)_{I}(\mathbb{R})=\operatorname{span}_{\mathbb{R}},\{L_{u,v}|u,v\in\mathcal{V}\}\oplus\mathbb{R}h\oplus\mathbb{R}$
$\mathcal{D}^{2,-1}(V)_{I}(\mathbb{R})=\{\psi^{Q}(\tau^{Q}(v))|v\in\mathcal{V}\}.$
Similar to proposition 8.2.1, we can show that $\mathfrak{s}(\mathbb{R})\simeq\mathfrak{o}(\mathcal{V}\oplus\mathcal{V}_{1,1})$, where $\mathcal{V}_{1,1}$ is a two dimensional Lorentzian quadratic space.
8.2.3 The isomorphism $\mathfrak{s}(\mathbb{R})\simeq\mathfrak{s}^{*}(\mathbb{R})$
In Section 2.2.6 we introduced the real from $\mathfrak{s}(\mathbb{R})$ (that was further discussed in in Section 8.2.2). The main real from of $\mathfrak{s}$ that is used in this paper is $\mathfrak{s}(\mathbb{R})^{*}$ that was defined in Section 5.2.2.
In this section we build an explicit isomorphism between $\mathfrak{s}(\mathbb{R})$ and $\mathfrak{s}^{*}(\mathbb{R})$.
We define a linear invertible transformation $\mathcal{F}$ of $\mathcal{D}(C)$ by $\mathcal{F}(d)=\exp(\operatorname{ad_{i\frac{\pi}{4}h}})$. On $\mathcal{D}^{k,\ell}(C)$, $\mathcal{F}$ is given by multiplication by $e^{i\frac{\ell\pi}{2}}$.
Proposition.
The restriction $\mathcal{F}|_{\mathfrak{s}^{*}(\mathbb{R})}$ is an isomorphism of real Lie algebras from $\mathfrak{s}^{*}(\mathbb{R})$ onto $\mathfrak{s}(\mathbb{R})$. Moreover, it is the unique $O(q)\times\mathbb{R}^{\times}$-equivariant isomorphism of real Lie algebras from $\mathfrak{s}^{*}(\mathbb{R})$ onto $\mathfrak{s}(\mathbb{R})$ that on $\mathcal{D}^{0,1}(C)\cap\mathfrak{s}^{*}(\mathbb{R})$ is given by multiplication by $i$.
This immediately follows from the commutation relations in Appendix 8.1.
8.3 A Lie-theoretic description of $\Psi^{Q}$
In Section 2.2.3 we introduced the isomorphism $\Psi^{Q}:\mathcal{D}^{0,1}(C)\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{D}^{2,-1}(C)$. In this section we give a Lie-theoretic characterization of it using the Jacobson-Morozov theorem.
Recall that $h=2\chi+n-2\in\mathfrak{s}$. Starting with a $w\in V^{*}$, such that $Q^{*}(w)\neq 0$, the triple $\{e_{w},h,f_{w}:=Q^{*}(w)^{-1}\psi^{Q}(w)\}$ is an $\mathfrak{sl}_{2}$-triple in $\mathfrak{s}$ satisfying the standard commutation relations
$h,e_{w}]=2e_{w},\quad[h,f_{w}]=-2f_{w},\quad[e_{w},f_{w}]=h.$
The Jacobson-Morozov theorem (see e.g., [MR1920389, Ch. X]) for the finite-dimensional simple Lie algebra $\mathfrak{s}$ implies that for $w$ as above there is a unique $f_{w}\in\mathfrak{s}$ such that $\{e_{w},h,f_{w}\}$ is an $\mathfrak{sl}_{2}$-triple in $\mathfrak{s}$ satisfying the standard commutation relations. This uniquely determines the morphism $\psi^{Q}$.
Remark.
The Lie algebra generated by the above triple $\{e_{w},h,f_{w}\}$ is $\ell_{w}\simeq\mathfrak{sl}_{2}(\mathbb{C})$ that was introduced in Section 3.0.2 as a part of a dual pair.
We remark that the one-dimensional subspace spanned by $h$ is given by the intersection of all such $\ell_{w}$. Also note that $h$ is the unique element in $\mathbb{C}h\subset\ell_{w}$ satisfying $h,e_{w}]=2e_{w}$.
8.4 On the reductive dual pair symmetry of $\mathfrak{s}$
Here we give a proof of Proposition 3.0.2.
Proof.
Using the explicit commutation relations in $\mathfrak{s}$}, which were given in Appendix 8.1, it can be shown that $\ell_{w}=\operatorname{span}_{\mathbb{C}}(\Psi^{Q}\left({w}\right),{w},h)\cong\mathfrak{sl}_{2}(\mathbb{C})$ and $\mathfrak{k}_{w}=\operatorname{span}_{\mathbb{C}}\{L_{u,v}|u,v\in V_{w}^{\perp}\}\simeq\mathfrak{o}(n-1,\mathbb{C})$, where $V_{w}^{\perp}:=\{v\in V|B(w,v)=0\}$. These facts together with the explicit commutation relations finish the proof of 1. The proof of 2. is similar. ∎
8.5 Concrete realization of $(\boldsymbol{H}^{\infty}_{\pm})_{\ell}$
As explained in section 7.2.3, the unitary representation $\pi_{\ell}$ of $SL_{2}(\mathbb{R})\times SO(3)$ on $\boldsymbol{H}_{\ell}$ can be realized on $L^{2}\left(\mathbb{R}\setminus\{0\},\frac{1}{2}|t|dt\right)\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}$. It follows from Sections 7.2.3, and 7.2.5 that in this realization
$(\boldsymbol{H}_{+})_{\ell}=L^{2}\left((0,\infty),\frac{1}{2}|t|dt\right)\otimes_{\mathbb{C}}\mathcal{Y}^{\ell},$
and
$(\boldsymbol{H}_{-})_{\ell}=L^{2}\left((-\infty,0),\frac{1}{2}|t|dt\right)\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}.$
We denote the corresponding irreducible subrepresentations of $SL_{2}(\mathbb{R})\times SO(3)$ by $(\pi_{\ell,+},(\boldsymbol{H}_{+})_{\ell})$ and $(\pi_{\ell,-},(\boldsymbol{H}_{-})_{\ell})$.
We recall that the Schwartz space $\mathcal{S}([0,\infty))$ of rapidly decreasing functions on $0,\infty)$ is given by the space of all complex-valued functions on $0,\infty)$ that are smooth on $(0,\infty)$, smooth from the right at 0, and rapidly decaying at infinity; explicitly, for any $m,n\in\mathbb{N}_{0}$, $\sup\{|t^{m}f^{(m)}(t)|:t\in[0,\infty)\}<\infty$.
Similarly, the Schwartz space $\mathcal{S}((-\infty,0])$ of rapidly decreasing functions on $(-\infty,0]$ is defined.
Proposition.
For every $\ell\in\mathbb{N}_{0}$,
  1. 1.
    The smooth subspace $(\boldsymbol{H}^{\infty}_{+})_{\ell}$ of $(\boldsymbol{H}_{+})_{\ell}$ is $t^{\ell}\mathcal{S}([0,\infty))\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}$.
  2. 2.
    The smooth subspace $(\boldsymbol{H}^{\infty}_{-})_{\ell}$ of $(\boldsymbol{H}_{-})_{\ell}$ is $t^{\ell}\mathcal{S}((-\infty,0])\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}$.
Proof.
The proof immediately follows form the description of the smooth subspace of the discrete series representations of $SL_{2}(\mathbb{R})$ in the Kirillov model as given in [Baruch]. ∎
8.6 Spectrum of the Schrödinger family is always real
Here we show that the spectrum of the Schrödinger family in any admissible smooth Fréchet representation of moderate growth is a subset of the real numbers.
Proposition.
Let $(\pi,SL_{2}(\mathbb{R}),{W})$ be an admissible smooth Fréchet representation of moderate growth. Then for any $\kappa>0$
$\operatorname{Spec}_{\boldsymbol{W}}\{S^{Ab}_{\kappa}\}\subseteq\mathbb{R}.$
Proof.
Lemma 7.1.2 shows that it is enough to prove the proposition for $\kappa=1$. We can further assume that $W$ is irreducible. By Casselman submodule theorem we can assume that $W$ is a quotient of a principal series representation realized on the space of $\nu$-homogeneous smooth functions on the punctured plane $\mathbb{R}^{2}_{0}:=\mathbb{R}^{2}\setminus\{0\}$ for some $\nu\in\mathbb{C}$. Restriction to the circle $S^{1}\subset\mathbb{R}^{2}_{0}$ gives an isomorphic representation on the space $C^{\infty}(S^{1})$ of smooth functions on the circle.
In this representation the Schrödinger family acts via
$\pi^{\nu}(S^{Ab}_{1}({\lambda}))=\left(1+i\nu(\lambda+1)s(\theta)\right)+ic_{\lambda}(\theta)\frac{d}{d\theta},$
where
$s(\theta):=\sin(\theta)\cos(\theta),\quad c_{\lambda}(\theta):=\cos^{2}(\theta)-\lambda\sin^{2}(\theta).$
The dual space of $C^{\infty}(S^{1})$ is the space of distributions on $S^{1}$. The adjoint operator
$\pi^{\nu}(S^{Ab}_{1}({\lambda}))^{*}:D^{\prime}(S^{1})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow D^{\prime}(S^{1})$
is given by
$\pi^{\nu}(S^{Ab}_{1}({\lambda}))^{*}=$ $1+i(\nu+2)(\lambda+1)s(\theta)-ic_{\lambda}(\theta)\frac{d}{d\theta}$
For $\lambda\in\mathbb{C}\setminus\mathbb{R}$ (in fact this is true for $\lambda\in\mathbb{C}\setminus[0,\infty)$), the leading coefficient $c_{\lambda}(\theta)\neq 0$ never vanishes. Hence for $\lambda\in\mathbb{C}\setminus\mathbb{R}$, any distribution in $\operatorname{Ker}(\pi^{\nu}(S^{Ab}_{1}({\lambda}))^{*})$ must be a smooth function on the circle which locally must be of the form
$C(c_{\lambda}(\theta))^{-(\nu+2)/2}\exp\left\{-i\int_{\theta_{0}}^{\theta}\frac{dt}{c_{\lambda}(t)}\right\},$
for some complex constant $C$ and some real constant $\theta_{0}$. For existence of a global nonzero solution we must demand that the monodromy $M$, when varying $\theta$ along $0,2\pi]$, is equal to $1$.
The monodromy can be calculated explicitly, it is independent of $\nu$ and is given by
$M=\exp\left\{-i\int_{0}^{2\pi}\frac{d\theta}{c_{\lambda}(\theta)}\right\}=\exp\left\{-i\int_{0}^{2\pi}\frac{c_{\overline{\lambda}}(\theta)d\theta}{|c_{\lambda}(\theta)|^{2}}\right\}.$
Hence
$|M|=\exp\left\{\int_{0}^{2\pi}\frac{\operatorname{Im}(c_{\overline{\lambda}}(\theta))d\theta}{|c_{\lambda}(\theta)|^{2}}\right\}=\exp\left\{-\operatorname{Im}(\lambda)\int_{0}^{2\pi}\frac{\sin^{2}(\theta)d\theta}{|c_{\lambda}(\theta)|^{2}}\right\}.$
Since we assume that $\lambda\notin\mathbb{R}$ and since the integral is positive $|M|\neq 1$. ∎
We remark that similar but more detailed calculations of the monodromy show that
$\operatorname{Spec}_{\boldsymbol{W}}\{S^{Ab}_{\kappa}\}\subseteq\left\{-\frac{\kappa^{2}}{n^{2}}|n\in\mathbb{N}\right\}\cup[0,\infty).$
8.7 The self-adjoint property of the subspace of smooth vectors
In the Section we prove Claim 5.1.4.
Let $G$ be a real Lie group, $\mathfrak{g}$ the complexification of the Lie algebra $\operatorname{Lie}(G)$, and $\mathcal{U}(\mathfrak{g})$ its universal enveloping algebra.
Let $(\pi,G,\boldsymbol{H})$ be a unitary representation of $G$ on a Hilbert space $\boldsymbol{H}$. Denote by $\boldsymbol{H}^{\infty}$ the subspace of $\boldsymbol{H}$ consisting of smooth vectors for $\pi$.
Claim.
The natural action $\pi$ of the universal enveloping algebra $\mathcal{U}(\mathfrak{g})$ on the space $\boldsymbol{H}^{\infty}$ is self-adjoint.
Before proving the claim we introduce relevant Sobolev spaces and other tools that will be used in the proof.
We denote by $\mathcal{U}^{n}(\mathfrak{g})$ be the $n$-th piece in the PBW filtration of $\mathcal{U}(\mathfrak{g})$ consisting of elements of degree $\leq n$.
We fix a *-invariant Hermitian scalar product $B_{n}$ on $\mathcal{U}^{n}(\mathfrak{g})$ and denote the corresponding Hermitian scalar product on $T^{n}:=\mathcal{U}^{n}(\mathfrak{g})\otimes_{\mathbb{C}}\boldsymbol{H}$ by $\sigma_{n}$.
For $n\in\mathbb{Z}_{\geq 0}$, the $n$-th Sobolev norm of $v\in\boldsymbol{H}^{\infty}$ is defined as the operator norm of the morphism
$M_{v}:(\mathcal{U}^{n}(\mathfrak{g}),B_{n})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow(\boldsymbol{H},\langle\_,\_\rangle)$
$X\longmapsto\pi^{\infty}(X)v.$
The Fréchet topology on $\boldsymbol{H}^{\infty}$ is induced by this sequence of Sobolev norms.
The adjoint action of $G$ on $\mathcal{U}^{n}(\mathfrak{g})$ is continuous and we obtain a continuous representation of $G$ on $T^{n}$. The smooth subspace of $T^{n}$ is clearly given by $T^{n,\infty}=\mathcal{U}^{n}(\mathfrak{g})\otimes_{\mathbb{C}}\boldsymbol{H}^{\infty}$.
The action $\pi$ defines a continuous linear map $\pi^{n}:T^{n,\infty}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}$ via
$X\otimes w\longmapsto\pi(X)w.$
Hence any vector $v\in\boldsymbol{H}$ defines a linear functional $\xi_{v}$ on $T^{n,\infty}$ (in fact on all $\mathcal{U}(\mathfrak{g})\otimes_{\mathbb{C}}\boldsymbol{H}^{\infty}$) via
$\xi_{v}(X\otimes w)=\langle\pi(X)w,v\rangle.$
Definition.
The $n$-th Sobolev space $\boldsymbol{H}^{n}\subset\boldsymbol{H}$ consists of all vectors $v\in\boldsymbol{H}$ such that the linear functional $\xi_{v}$ is bounded with respect to norm associated with $\sigma_{n}$. We denote the norm of the functional $\xi_{v}$ by $\tau_{n}(v)$. The map $\tau_{n}:\boldsymbol{H}^{n}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathbb{R}_{\geq 0}$ is a norm on $\boldsymbol{H}^{n}$.
In the proof of the claim we shall use the following lemma that can be easily verified.
Lemma.
  1. 1.
    The norm $\tau_{n}$ on $\boldsymbol{H}^{n}$ is induced by an Hermitian inner product $\langle\_,\_\rangle_{n}$. With respect to $\langle\_,\_\rangle_{n}$, $\boldsymbol{H}^{n}$ is a Hilbert space.
  2. 2.
    Each $\boldsymbol{H}^{n}$ is stable under $\pi(G)$, and this action of $G$ is continuous.
  3. 3.
    $\boldsymbol{H}^{\infty}\subset\boldsymbol{H}^{n}$ and restriction of $\tau_{n}$ to $\boldsymbol{H}^{\infty}$ coincides with the the $n$-th Sobolev norm on $\boldsymbol{H}^{\infty}$ that was introduced before.
Proof of the claim.
From the definition of the adjoint it is clear that $(\boldsymbol{H}^{\infty})^{\dagger}=\cap_{n}\boldsymbol{H}^{n}$. Note that for $n>m$ the inclusion $\boldsymbol{H}^{n}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}^{m}$ is a continuous $G$-equivariant map. Hence it is enough to show that $\boldsymbol{H}^{\infty}=\cap_{n}\boldsymbol{H}^{n}$. Clearly the inclusion $\boldsymbol{H}^{\infty}\subseteq\cap_{n}\boldsymbol{H}^{n}$ is a topological embedding. We will prove the required equality using the fact the $\boldsymbol{H}^{\infty}$ is complete with respect to the system of Sobolev norms $\tau_{n}$. This follows from [BK2014, Lemma. 2.15] (also see [MR84713]).
Let $v\in\cap_{n}\boldsymbol{H}^{n}\subset\boldsymbol{H}$. For any smooth measure $\mu$ with compact support on $G$ we have $\pi(\mu)v\in\boldsymbol{H}^{\infty}\subseteq\cap_{k}\boldsymbol{H}^{k}$. Using continuity of the actions of $G$ on the Sobolev spaces, for any $n\in\mathbb{Z}_{\geq 0}$ we can find a smooth measure $\mu_{n}$ on $G$, supported in a small neighborhood of the identity in $G$, such that $\tau_{m}(v-\pi(\mu_{n})v)<2^{-n}$ for all $m\leq n$.
The above inequality shows that the sequence of vectors $v_{n}:=\pi(\mu_{n})v$ is a Cauchy sequence in $\boldsymbol{H}^{\infty}$ with respect to all norms $\tau_{n}$. Using completeness of $\boldsymbol{H}^{\infty}$, this sequence has a limit in $\boldsymbol{H}^{\infty}$. The unique limit is clearly $v$, hence $v\in\boldsymbol{H}^{\infty}$. ∎

### 14 Apr 2026 v2

@@
 1.1 Overview
-The hydrogen atom system is a fundamental well-studied example of a quantum-mechanical system which played a central role in the development of quantum mechanics, see e.g., [Wybourne, LandauLifshitz, MR2156403, Wulfman].
+The hydrogen atom system is a fundamental well-studied example of a quantum-mechanical system which played a central role in the development of quantum mechanics, see e.g., [undeff, undefs, undefad, undefai].
 The common model for the spinless non-relativistic system is given by the Hilbert space $L^{2}(\mathbb{R}^{3})$, together with the densely defined (normalized) Schrödinger operator
@@
 These solution spaces lead to several layers of symmetries in which the largest symmetry group is $O(4,2)$.
-In physics, it is called the dynamical symmetry e.g., see [sym12081323].
+In physics, it is called the dynamical symmetry e.g., see [undefw].
 The model we just described has several flaws.
@@
 We noticed two a priori unrelated situations in which there is an expected symmetry of $SO(n)$ but in practice the symmetry is of a larger orthogonal group.
-The first instance is the usual model of the hydrogen atom system, in which there is a clear symmetry of $SO(3)$ and a larger hidden symmetry [^1] In fact there are different hidden symmetry groups within the dynamical symmetry that are fibers of a nice family, See [MR4460278, MR3827131].
+The first instance is the usual model of the hydrogen atom system, in which there is a clear symmetry of $SO(3)$ and a larger hidden symmetry [^1] In fact there are different hidden symmetry groups within the dynamical symmetry that are fibers of a nice family, See [undefh, undefaf].
 of $O(4,2)$ that is called the dynamical symmetry.
@@
 This was the starting point in our search for a new model for the hydrogen atom system which eventually lead to the results presented in this paper.
+Recently we discovered works by Meng [undefx, undefy] that introduce a model close to our model. Meng considers a more general case that includes the case of the Lorentz group acting on a cone.
+Meng starts from a simple Euclidean Jordan algebra $J$ and builds a Lie algebra $\mathfrak{co}$, the conformal Lie algebra of $J$, using the Kantor–Koecher–Tits construction.
+Then Meng constructs a Kepler cone $\mathcal{C}$ associated to $J$ and considers the Hilbert space $\mathcal{H}=L^{2}(\mathcal{C})$.
+He observes that the Lie algebra $\mathfrak{co}$ acts on functions on $\mathcal{C}$ via differential operators.
+Then he defines a Schrödinger operator in terms of the Lie algebra $\mathfrak{co}$ and computes its discrete spectrum.
+In the case of a Jordan algebra $J$ corresponding to Lorentzian space the Kepler cone $\mathcal{C}$ constructed by Meng coincides with our cone $C_{+}(\mathbb{R})$, the Hilbert space $\mathcal{H}=L^{2}(\mathcal{C})$ is our space $\boldsymbol{H}_{+}=L^{2}(C_{+}(\mathbb{R}))$ and his Lie algebra $\mathfrak{co}$ is isomorphic to (a real form of) our quantum dynamical symmetry $\mathfrak{s}$ (see subsection 2.2.2).
+Our approach is much more algebraic.
+We mostly work with the algebra $\mathcal{D}(C)$ of algebraic differential operators on $C$; the Lie algebra $\mathfrak{s}$ arises as a convenient generating subspace of $\mathcal{D}(C)$ and our Schrödinger family lives inside $\mathcal{D}(C)$.
+Also, one of the main ingredients of our model is the Schwartz space $\boldsymbol{H}^{\infty}\subset\boldsymbol{H}$ and a self-adjoint action of the algebra $\mathcal{D}(C)$ on the Schwartz space.
+The Schwartz space implicitly includes boundary conditions. This allows us to investigate the continuous spectrum in addition to the discrete spectrum.
 1.2 Structure of the paper and main results
@@
 When $X$ is smooth, $\mathcal{D}(X)$ is generated as an algebra by $\mathcal{O}_{X}$ and $\operatorname{Der}(\mathcal{O}_{X})=\{D\in\mathcal{D}^{1}(X)|D(1)=0\}$.
-In general, beyond the smooth case, $\mathcal{D}(X)$ may not even be Noetherian, e.g., see the case of the cubic cone [cubiccone].
+In general, beyond the smooth case, $\mathcal{D}(X)$ may not even be Noetherian, e.g., see the case of the cubic cone [undefa].
 Consider the case in which the algebraic variety $X$ is a finite-dimensional complex vector space $V$.
@@
 $\mathcal{D}(V)_{I}=\{D\in\mathcal{D}(V)|D(I)\subset I\}.$
-The following result which follows from Proposition 1.6 of [SSP], explains how a differential operator on the cone can be described by differential operators on the ambient vector space.
+The following result which follows from Proposition 1.6 of [undefae], explains how a differential operator on the cone can be described by differential operators on the ambient vector space.
 Proposition.
@@
 It can be verified that $\tilde{\mathfrak{s}}=\mathfrak{s}\oplus\mathbb{C}$.
-Theorem (Compare with Prop. 3.8 and Cor. 4.4 of [LevasseurAndStafford2017]).
+Theorem (Compare with Prop. 3.8 and Cor. 4.4 of [undefu]).
 The complex Lie algebra $\mathfrak{s}$ is isomorphic to $\mathfrak{so}(n+2,\mathbb{C})$ and it generates $\mathcal{D}(C)$ as a unital complex algebra.
-For an explicit embedding of $\mathfrak{so}(n+2,\mathbb{C})$ in $\mathcal{D}(C)$ see section 9 of [LevasseurAndStafford2017] and Theorem 8.2.1.
-Further details can be found in [Levasseur1986, LevasseurSmithStafford1988, Goncharov1982] and section 1.1 of [KobayashiMano2011].
+For an explicit embedding of $\mathfrak{so}(n+2,\mathbb{C})$ in $\mathcal{D}(C)$ see section 9 of [undefu] and Theorem 8.2.1.
+Further details can be found in [undefv, undeft, undefg] and section 1.1 of [undefp].
 Remark.
@@
 By direct calculations we see that $\Psi^{Q}(V^{*})\subseteq\mathcal{D}^{2,-1}(V)_{I}$ and for every $w\in V^{*}$, $\Phi^{Q}\circ\Psi^{Q}(w)=(2-n)w$.
-In addition it is easy to show (and can be deduced from [Goncharov1982]) that $\operatorname{dim}(\mathcal{D}^{2,-1}(V)_{I})\leq\operatorname{dim}(V)$.
+In addition it is easy to show (and can be deduced from [undefg]) that $\operatorname{dim}(\mathcal{D}^{2,-1}(V)_{I})\leq\operatorname{dim}(V)$.
 ∎
@@
 $\mathfrak{s}\cong V\oplus\mathfrak{o}(Q)\oplus\mathbb{C}\oplus V^{*}.$
-We remark that a description of $\mathcal{D}(C)$ is essentially given in [Goncharov1982]. We will also give an explicit description in the following section.
+We remark that a description of $\mathcal{D}(C)$ is essentially given in [undefg]. We will also give an explicit description in the following section.
 2.2.5 A spanning set for $\tilde{\mathfrak{s}}$
@@
 Throughout this paragraph, we let $(\mathcal{V},q,w)$ be an $n$-dimensional PLQS with $n>2$.
-Howe introduced the notion of a reductive dual pair, see e.g., [MR789290].
+Howe introduced the notion of a reductive dual pair, see e.g., [undefj].
 Given a reductive Lie algebra $\mathfrak{g}$ and a pair of reductive Lie subalgebras $(\mathfrak{k},\mathfrak{l})$ of $\mathfrak{g}$, the pair is called a reductive dual pair if $\mathfrak{k}$ is equal to the centralizer of $\mathfrak{l}$ in $\mathfrak{g}$ and $\mathfrak{l}$ is equal to the centralizer of $\mathfrak{k}$ in $\mathfrak{g}$.
@@
 5.1 Action of a *-algebra on a Hilbert space
-In this section we recall the notion of a partial action of a *-algebra on a Hilbert space and discuss some of its properties. We essentially follow [MR283580] (also see [MR333743, MR3722559]).
+In this section we recall the notion of a partial action of a *-algebra on a Hilbert space and discuss some of its properties. We essentially follow [undefaa] (also see [undefab, undefz]).
 5.1.1 A partial action of a *-algebra
@@
 5.1.3 Symmetric, self-adjoint and essentially self-adjoint actions
-Definition (Compare with Sec. II &\& IV of [MR283580]).
+Definition (Compare with Sec. II &\& IV of [undefaa]).
   1. 1.
@@
 The natural action $\pi^{\infty}$ of $\mathcal{U}(\mathfrak{g})$ on $\boldsymbol{H}^{\infty}$ is a dense self-adjoint action.
-We will prove this claim in Appendix 8.7. Further information on the underlying smooth space of a representation of a Lie group can be found in [BK2014].
+We will prove this claim in Appendix 8.7. Further information on the underlying smooth space of a representation of a Lie group can be found in [undefb].
 5.2 The canonical symmetric action of $\mathcal{D}(C)$ on $\boldsymbol{H}$
@@
 The representation of $L$ on $\boldsymbol{H}$ extends to a representation $\pi=\pi_{n}$ of $P$, by defining $\pi(n)$ for $n\in N$ to be the operator of multiplication by the function on $C_{0}(\mathbb{R})$ given by $v\longmapsto e^{iB(nv_{2},v)}$.
-By Mackey theory this unitary representation of $P$ is irreducible, also see [KOBAYASHI2003551, Prop. 3.3].
-The extension of the representation to $G$: The extension of $\pi_{n}=\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H})\right)$ to a representation of $G$ is based on the fact that $G$ has a distinguished representation $\widetilde{\pi}$ of minimal Gelfand-Kirillov dimension, namely the local theta lift of the trivial representation of $SL_{2}(\mathbb{R})$, see [MR1457244].
-Further info on the theta correspondence can be found in [MR546602, MR985172, MR986027].
+By Mackey theory this unitary representation of $P$ is irreducible, also see [undefr, Prop. 3.3].
+The extension of the representation to $G$: The extension of $\pi_{n}=\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H})\right)$ to a representation of $G$ is based on the fact that $G$ has a distinguished representation $\widetilde{\pi}$ of minimal Gelfand-Kirillov dimension, namely the local theta lift of the trivial representation of $SL_{2}(\mathbb{R})$, see [undefaj].
+Further info on the theta correspondence can be found in [undefi, undefl, undefk].
 Theorem.
@@
 Uniqueness follows from irreducibility of $\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H}\right)$ together with Schur’s lemma.
-The hard part of the proof is to show the existence of such an extension. This is proven in [KOBAYASHI2003551, Thm. 3]. ∎
+The hard part of the proof is to show the existence of such an extension. This is proven in [undefr, Thm. 3]. ∎
 Remark.
@@
 $\pi(-I)=(-1)^{(n-2)/2}.$
-2. For $n>4$ the representation $\widetilde{\pi}_{n}$ is the unique minimal representation of $G$, see [KobayashiMano2011].
+2. For $n>4$ the representation $\widetilde{\pi}_{n}$ is the unique minimal representation of $G$, see [undefp].
 3.
@@
 Other important realizations of $\widetilde{\pi}$ exist.
-Several of them are constructed in [KobayashiMano2011].
-Also see [MR2020550, MR1108044, MR1172839].
+Several of them are constructed in [undefp].
+Also see [undefq, undefc, undefm].
 5.3.2 The smooth subspace of $\pi$ as a self-adjoint extension
@@
 Proof.
-The first part of the proposition is proven in [KobayashiMano2011, Sec.
+The first part of the proposition is proven in [undefp, Sec.
 2.5].
 From the explicit formulas by which $\operatorname{Lie}(G)$ acts on $\boldsymbol{H}^{\infty}$ via differential operators on $C_{0}(\mathbb{R})$ e.g.
-see [KobayashiMano2011, Sec.
+see [undefp, Sec.
 2.3], it is evident that the self-adjoint action $\pi^{\infty}:\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{H}^{\infty})$ where $\mathfrak{g}$ is the complexification of the Lie algebra $\operatorname{Lie}(G)$ (as in Subsection 5.1.4 ) factors through a quotient surjective map $j:\mathcal{U}(\mathfrak{g})\twoheadrightarrow\mathcal{D}(C)$.
@@
 2.
-The decomposition $\boldsymbol{H}=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}$ is stable under the connected component of the identity $G_{0}\subset G$, see [KobayashiMano2011, p.19].
+The decomposition $\boldsymbol{H}=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}$ is stable under the connected component of the identity $G_{0}\subset G$, see [undefp, p.19].
 This implies that $\boldsymbol{H}^{\infty}=\boldsymbol{H}^{\infty}_{+}\oplus\boldsymbol{H}^{\infty}_{-}$, where $\boldsymbol{H}^{\infty}_{\pm}:=\boldsymbol{H}_{\pm}\cap\boldsymbol{H}^{\infty}$ and moreover $\boldsymbol{H}^{\infty}_{\pm}$ are stable under $\rho$.
@@
 7.1.1 The spectrum of a family of operators
-In functional analysis, given a continuous operator $T$ on a dense linear subspace $\boldsymbol{W}$ of a topological vector space $\boldsymbol{H}$ (typically $\boldsymbol{H}$ is a Hilbert space) the spectrum of $T$ is defined as the collection of all $\lambda\in\mathbb{C}$ for which the operator $T-\lambda:\boldsymbol{W}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}$ does not have a continuous inverse, see e.g., Section VI3 of [reed].
+In functional analysis, given a continuous operator $T$ on a dense linear subspace $\boldsymbol{W}$ of a topological vector space $\boldsymbol{H}$ (typically $\boldsymbol{H}$ is a Hilbert space) the spectrum of $T$ is defined as the collection of all $\lambda\in\mathbb{C}$ for which the operator $T-\lambda:\boldsymbol{W}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}$ does not have a continuous inverse, see e.g., Section VI3 of [undefac].
 In physics, given an operator $T$ as above, one is usually interested in a spectral decomposition of $\boldsymbol{H}$ with respect to $T$.
@@
 $(\boldsymbol{H}_{\pm})_{\ell}\cong D^{\pm}_{2\ell+2}\otimes_{\mathbb{C}}\mathcal{Y}^{\ell}$
-This follows from Sections 1.3 and 3.1 of [MR2401813].
+This follows from Sections 1.3 and 3.1 of [undefo].
 Remark.
@@
 However, in this paper we solve this problem working with these representations of $SL_{2}(\mathbb{R})$ as they concretely appear inside the Hilbert space $\boldsymbol{H}$ of functions on the cone.
-By Casselman-Wallach globalization theorem [Wallach3, Casselman, Wallach2], any two admissible smooth moderate growth Fréchet representations of the same reductive Lie group that are infinitesimally equivalent must be isomorphic as topological representations.
+By Casselman-Wallach globalization theorem [undefag, undefe, undefah], any two admissible smooth moderate growth Fréchet representations of the same reductive Lie group that are infinitesimally equivalent must be isomorphic as topological representations.
 The topological vector space $\boldsymbol{H}^{\infty}$ is admissible smooth moderate growth Fréchet representation of $G\simeq O(4,2)$.
@@
 $h,e_{w}]=2e_{w},\quad[h,f_{w}]=-2f_{w},\quad[e_{w},f_{w}]=h.$
-The Jacobson-Morozov theorem (see e.g., [MR1920389, Ch.
+The Jacobson-Morozov theorem (see e.g., [undefn, Ch.
 X]) for the finite-dimensional simple Lie algebra $\mathfrak{s}$ implies that for $w$ as above there is a unique $f_{w}\in\mathfrak{s}$ such that $\{e_{w},h,f_{w}\}$ is an $\mathfrak{sl}_{2}$-triple in $\mathfrak{s}$ satisfying the standard commutation relations.
@@
 Proof.
-The proof immediately follows form the description of the smooth subspace of the discrete series representations of $SL_{2}(\mathbb{R})$ in the Kirillov model as given in [Baruch]. ∎
+The proof immediately follows form the description of the smooth subspace of the discrete series representations of $SL_{2}(\mathbb{R})$ in the Kirillov model as given in [undef]. ∎
 8.6 Spectrum of the Schrödinger family is always real
@@
 We will prove the required equality using the fact the $\boldsymbol{H}^{\infty}$ is complete with respect to the system of Sobolev norms $\tau_{n}$.
-This follows from [BK2014, Lemma.
-2.15] (also see [MR84713]).
+This follows from [undefb, Lemma.
+2.15] (also see [undefd]).
 Let $v\in\cap_{n}\boldsymbol{H}^{n}\subset\boldsymbol{H}$.